### v3

# Quantum Advantage via Efficient Post-processing on Qudit Shadow Tomography

Abstract
The calculation of $\text{tr}(AB)$ is essential in fields like quantum science and artificial intelligence, but the classical computational complexity is $O(d^{2})$ when $A$ and $B$ are $d$-dimensional matrices. Moreover, storing $A$ and $B$ requires $O(d^{2})$ memory, which poses additional challenges for exponential high-dimensional systems. We propose a quantum approach through a qudit shadow tomography framework to exponentially reduce both the computational and storage complexity to $O(\text{poly}(\log d))$ for a broad class of matrices $A$ and for bounded-norm Hermitian matrices $B$ with known $\text{tr}(B)$. Compared to shadow tomography via random Clifford measurements, our method reduces the computational complexity of post-processing per measurement from an exponential worst-case scenario to a constant, and it is applicable across arbitrary dimensions $d$. This advancement opens new pathways for efficient high-dimensional data analysis and complex system modeling.
Introduction—Quantum advantage represents a transformative leap in computational power, breaking classical limits across various fields. From the exponential query complexity reduction in the Deutsch-Jozsa algorithm [1] to Shor’s algorithm for breaking RSA encryption [2] and the HHL algorithm’s ability to solve linear equations exponentially faster [3], quantum algorithms continue to reshape what is computationally feasible. Recent advancements, such as random circuit sampling on NISQ devices [4, 5], underscore the potential for groundbreaking applications. With increasing demands for high-dimensional data processing, particularly in artificial intelligence, advancing quantum advantage holds transformative potential.
Efficiently computing $\text{tr}(AB)$ is a central task across quantum science, machine learning, optimization, and data analysis, especially for matrix operations, quantum measurements, kernel methods, and optimization problems [6, 7]. However, as dimension $d$ grows exponentially (e.g., $d=2^{n}$), both storage and computation of traces become computationally demanding. In contrast, an $n$-qubit quantum state can represent a $2^{n}$-dimensional density matrix, which can exponentially reduce storage requirements in quantum simulations [8].
This work targets a specific subset of matrices $A$, particularly Hermitian, positive semi-definite matrices that represent the density matrix of a quantum state (with trace 1), while $B$ is a Hermitian matrix with a given bounded norm. In classical computation, both calculating $\text{tr}(AB)$ and storing matrices $A$ scale as $O(d^{2})$, requiring exponential resources as $d$ scales exponentially. We aim to overcome this limitation by leveraging quantum advantage to achieve exponential speedup in both computation and storage. Specifically, we target the computation of $\text{tr}(AB)$ with polynomial computational complexity, $O(\text{poly}(\log d))$, and arbitrary precision. This is especially impactful in high-dimensional settings, where classical methods become inefficient, thereby extending the reach of quantum computing to complex problems across various domains, as illustrated in Fig. 1.
Figure 1: Advantage seeking: In the absence of noise, the output of quantum computation is an $n$-qubit pure state $|x\rangle$, representing a $2^{n}$-dimensional vector. This state can serve as an intermediate step in quantum machine learning algorithms, such as solving $A\vec{x}=\vec{b}$ [3], and is commonly used as a sub-procedure in various quantum machine learning algorithms [9]. We assume that generating $|x\rangle$ through quantum computation is not exponentially slower than classical counterparts, though many quantum algorithms seek exponential speedup at this stage. All quantum states are represented by a density matrix $\rho$. Although $\rho$ has a limited lifetime, it can be stored as polynomial $n$ classical bits through well-designed random measurements. The main challenge is determining whether we can efficiently compute $\text{tr}(\rho O_{k})$ from these $n$ bits with polynomial-time complexity and arbitrary precision. Additionally, we explore whether this approach can be extended to a broader set of density matrices $\rho$ for bounded norm observables $O$.
We build on the shadow tomography framework [10], a crucial technique for efficiently predicting properties $\{\text{tr}(\rho O_{k})\}$ of an unknown quantum state $\rho$. Shadow tomography enables a significant quantum advantage in experimental learning, achieving exponential reductions in sample complexity compared to classical approaches [11]. Calculating $\mathop{\mathrm{tr}}(\rho O)$ for bounded norm Hermitian operators $O$, which are essential for evaluating global properties like fidelity estimation and entanglement detection, is crucial; however, shadow tomography based on random Clifford measurements and MUB measurements [12] encounter several challenges. Although sampling complexity is polynomial, post-processing for general observables lacking efficient representation can still be exponentially complex. Furthermore, in arbitrary $d$-level systems, neither the processing of the Clifford measurements nor the existence of $d+1$ MUBs is fully established [13]. On optical platforms, projecting onto random states rather than implementing circuits requiring calculation of all coefficients of Clifford and MUB states for $n$-qubit systems results in a time complexity of $O(2^{n}n^{3})$ [14]. These challenges highlight the need for refined approaches to improve the efficiency and applicability of shadow tomography across various quantum platforms.
In this work, we propose a novel qudit shadow tomography scheme for arbitrary $d$-dimensional systems, utilizing random projective measurements onto $2d$ Dense Dual Bases (DDBs). Unlike random Clifford measurements, which suffer from high post-processing complexity, our scheme ensures constant calculations per post-processing step, enabling polynomial-time computations as long as the sampling complexity remains polynomial. This method is particularly well-suited for optical systems, where projected states contain at most two non-zero amplitudes, making it both experimentally feasible and efficient.
Preliminaries and framework—Shadow tomography [10] is to predict an unknown quantum state’s properties efficiently. The process involves randomly performing operations from an informationally complete unitary ensemble $\mathcal{U}=\{U_{k}\}$. In each experiment, the state $\rho$ is evolved using $U_{k}$ with probability $p_{k}$, measured in the computational basis, and then the result $j$ is recorded. It corresponds to a quantum process
$\mathcal{M}(\rho)=\sum_{k}p_{k}\mathop{\mathrm{tr}}(U_{k}\rho U_{k}^{{\dagger}}|j\rangle\langle j|)U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}.$
Each experiment yields a single estimate of the unknown observable $O$ as
$\mathop{\mathrm{tr}}(O\mathcal{M}^{-1}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k})).$ (1)
Denote $\langle o_{k}\rangle$ as the estimation of $\mathop{\mathrm{tr}}(\rho O_{k})$. Since the quantum state collapses upon measurement, multiple copies of the unknown state should be prepared. Given exponentially large number of observables $\{O_{k}\}_{k=1}^{L}$, to achieve accuracy $\epsilon$ and confidence level $1-\sigma$, we can ensure
$P_{r}(|\langle o_{k}\rangle-\mathop{\mathrm{tr}}(\rho O_{k})|<\epsilon)\geq 1-\sigma$ (2)
with the following sample complexity:
$N=O\left(\frac{\log\frac{L}{\sigma}}{\epsilon^{2}}\right)\max_{1\leq i\leq L}\left\|O_{i}-\frac{\text{tr}(O_{i})}{d}I\right\|_{\mbox{shadow}}^{2},$ (3)
where $\|\cdot\|_{\mbox{shadow}}^{2}=\max_{\sigma:\text{ state }}\|\cdot\|_{\sigma}^{2}$ denotes the shadow norm associated with $\mathcal{U}$ and observable $O$. It represents the worst-case sample complexity, while $\|\cdot\|_{\sigma}^{2}$ is related to the sample complexity for state $\sigma$. Denote $O_{0}=O-\mathop{\mathrm{tr}}(O)I/d$, $\lVert O_{0}\rVert_{\sigma}^{2}$ is
$\mathbb{E}_{U\sim\mathcal{U}}\sum_{b\in\{0,1\}^{n}}\langle b|U\sigma U^{\dagger}|b\rangle\cdot\langle b|U\mathcal{M}^{-1}(O_{0})U^{\dagger}|b\rangle^{2}.$ (4)
If the ensemble $\mathcal{U}$ consists of all the $n$-qubit Clifford circuits, the reconstruction channel $\mathcal{M}^{-1}$ is given by
$\mathcal{M}^{-1}(\rho)=(2^{n}+1)U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}-\mathop{\mathrm{tr}}(\rho)I.$ (5)
The corresponding variance $\|\cdot\|_{\mbox{shadow}}\leq 3\mathop{\mathrm{tr}}^{2}(O_{0})$. Thus, polynomial sampling of the unknown state is sufficient for predicting bounded norm observables using random Clifford measurements.
Essentially, shadow tomography can be viewed as a quantum algorithm to approximate $\mathop{\mathrm{tr}}(\rho O)$ with arbitrary accuracy. For a bounded-norm observable $O$ and any density matrix $\rho$, the estimation using random Clifford measurements is given by:
$\mathop{\mathrm{tr}}(\rho O)\approx\sum_{k=1}^{\text{poly}(n)}\frac{(2^{n}+1)\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)-2^{n}\mathop{\mathrm{tr}}(O)}{\text{poly}(n)}.$ (6)
Here, $\text{poly}(n)$ corresponds to the sample complexity in Eq. (3). Statistical methods like the median of means are used to refine data reconstruction with partial data, but they do not reduce the overall sample complexity.
The overall computational complexity remains polynomial if each calculation of $\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)$ is constant or polynomial rather than exponential. The snapshots, or projected states, form the set
$\mathcal{S}=\{U_{k}^{{\dagger}}|j\rangle:U_{k}\in\mathcal{U},j=0,\cdots,d-1\}.$ (7)
However, the snapshots generated by random Clifford measurements are all stabilizer states. If the observable $O$ lacks an efficient representation in terms of states in $\mathcal{S}$, the calculation of $\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)$ in a single experiment can become exponential, posing challenges similar to those in the classical computation of $\mathop{\mathrm{tr}}(AB)$. This reflects a common issue related to the exponential post-processing complexity in shadow tomography.
To address this, we modify the collection of snapshots to ensure that the computational complexity of each calculation $\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)$ remains constant for any $O$.
Snapshots with dense dual basis states— We define the following states:
$\begin{cases}|\phi_{jk}^{\pm}\rangle=\frac{1}{\sqrt{2}}(|j\rangle\pm|k\rangle),\\ |\psi_{jk}^{\pm}\rangle=\frac{1}{\sqrt{2}}(|j\rangle\pm i|k\rangle).\end{cases}$ (8)
The new collection of snapshots comprises a total of $2d^{2}-d$ elements:
$\mathcal{S}_{\text{DDB}}=\{P_{t}=|t\rangle\langle t|,~{}~{}t=0,\ldots,d-1;\\ P_{jk}^{\pm}=|\phi_{jk}^{\pm}\rangle\langle\phi_{jk}^{\pm}|,Q_{jk}^{\pm}=|\psi_{jk}^{\pm}\rangle\langle\psi_{jk}^{\pm}|;~{}~{}0\leq j<k\leq d-1\}.$ (9)
These rank-1 projections are informationally complete and thus sufficient for shadow tomography.
An efficient algorithm with $O(\log d)$ time complexity has been developed to construct a unitary ensemble $\mathcal{U}=\{U_{j}\}_{j=1}^{f(d)}$ with the minimal number of elements required to span all states in Eq. (9) [15]. Each orthonormal basis $\{U_{j}|k\rangle:k=0,\dots,d-1\}$ is referred to as a Dense Dual Basis (DDB).
For even $d$, $f(d)=2d-1$. The first DDB corresponds to the computational basis, while the remaining bases are constructed in dual pairs, denoted as $\{|\phi_{jk}^{\pm}\rangle:(j,k)\in\mathbb{T}\}$ and $\{|\psi_{jk}^{\pm}\rangle:(j,k)\in\mathbb{T}\}$, where $\mathbb{T}$ is a partition of $\{0,1,\dots,d-1\}$ into distinct pairs with no repeated elements. For example, when $d=4$, one possible partition is $\mathbb{T}_{1}=\{(0,1),(2,3)\}$. A minimum of $d-1$ such partitions is needed to cover all pairs $(j,k)$ where $10\leq j<k\leq d-$ 1.
For odd $d$, $f(d)=2d$. In this case, the DDBs are similarly paired as $\{|\phi_{jk}^{\pm}\rangle,|l_{\mathbb{T}}\rangle:(j,k)\in\mathbb{T}\}$ and $\{|\psi_{jk}^{\pm}\rangle,|l_{\mathbb{T}}\rangle:(j,k)\in\mathbb{T}\}$, where $l_{\mathbb{T}}$ represents the single element in $\{0,1,\dots,d-1\}$ not included in $\mathbb{T}$. For example, with $d=5$ and $\mathbb{T}_{1}=\{(0,1),(2,3)\}$, the value of $l_{\mathbb{T}}$ is 4. A minimum of $d-1$ such partitions is also sufficient.
In $n$-qubit systems, each DDB circuit can be implemented using a Hadamard gate and a permutation gate, with a decomposition into $O(n^{4})$ elementary gates.
Compared with random Clifford measurements for shadow tomography, random DDB measurements offer several benefits. First, they are applicable across arbitrary dimensions. Second, in optical systems with random snapshot projections, the computational cost is significantly reduced. For instance, calculating a random stabilizer state requires $O(2^{n}n^{3})$ operations for dimension $d=2^{n}$ [14], whereas DDB states contain at most two nonzero amplitudes, greatly reducing computational demand. Third, the biased sampling process is more efficient; while the total number of stabilizer states scales as $O(2^{n^{2}})$, DDB states number only $O(2^{2n})$. Additionally, shadow tomography via random DDB measurements can also yield an explicit expression for the reconstruction channel.
Theorem 1 (Reconstruction channel).
For odd dimensions $d$, each unitary operation in the DDB ensemble $\{U_{j}\}_{j=1}^{2d}$ is sampled with probability $1/2d$. For even dimensions $d$, the ensemble $\{U_{j}\}_{j=1}^{2d-1}$ is sampled, with a probability of $2/2d$ assigned to $I$ (for computational basis states) and $1/2d$ to all other elements. The resulting quantum channel is
$\mathcal{M}(\rho)=\frac{1}{2d}\left[\rho+\text{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|\right].$ (10)
The corresponding reconstruction channel is
$\mathcal{M}^{-1}(\rho)=2d\left[\rho-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})P_{k}\right]-\frac{\text{tr}(\rho)}{d}I.$ (11)
The proof is detailed in Appendix A. In each single experiment, the calculation of $\mathop{\mathrm{tr}}\left[\mathcal{M}^{-1}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k})O\right]$ during post-processing becomes a constant. This efficiency arises because the term $U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}-\frac{d-1}{d}\sum_{k=0}^{d-1}\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}P_{k})P_{k}$ contains at most four nonzero elements, and $\mathop{\mathrm{tr}}(O)$ is known.
Theorem 2 (Performance guarantee).
In a $d$-dimensional Hilbert space, if we use random DDB measurements as shadow estimators for predicting the expectation value of a given observable $O$, worst-case variance for a quantum state $\sigma$ is bounded by
$\lVert O_{0}\rVert_{\text{shadow}}^{2}=\max_{\sigma:\text{ state}}\lVert O_{0}\rVert_{\sigma}^{2}\leq 2d\,\text{tr}(O_{0}^{2}),$ (12)
where $O_{0}=O-\mathop{\mathrm{tr}}(O)/d$. If the unknown state is sampled randomly according to the Haar measure, the average variance is bounded by
$\lVert O_{0}\rVert_{I/d}^{2}\leq 2\text{tr}(O_{0}^{2})$ (13)
The detailed proof is provided in Appendix B. For an arbitrary unknown quantum state $\rho$ (representing the Hermitian, positive semi-definite matrix $A$ with $\mathop{\mathrm{tr}}(A)=1$), the computational complexity of calculating $\mathop{\mathrm{tr}}(\rho O)$ using random DDB measurements is given by
$\mathcal{N}\leq O\left(\frac{\log\frac{L}{\sigma}}{\epsilon^{2}}\right)\times\lVert O_{0}\rVert_{\rho}^{2}\times\text{constant},$ (14)
where $\epsilon$ is the error relative to the exact value, and $1-\sigma$ denotes the confidence level.
When employing random DDB measurements, the exponential speedup in calculating $\mathop{\mathrm{tr}}(AB)$ (or $\mathop{\mathrm{tr}}(\rho O)$) can be achieved for all observables $O$ and specific states $\rho$ that satisfy $\lVert O_{0}\rVert_{\rho}^{2}=O(\text{poly}(\log d))$, given that the post-processing computational cost remains constant for any observable. However, in the worst-case scenario, only a quadratic speedup is achievable, as $\lVert O_{0}\rVert_{\text{shadow}}^{2}\leq 2d\mathop{\mathrm{tr}}(O^{2}_{0})$. In the average case, polynomial computational complexity is achievable when $\rho$ is randomly selected. Additionally, we will demonstrate that exponential speedup is attainable for all states and observables with specific intermediate processing.
In contrast, when utilizing random $n$-qubit Clifford measurements, the shadow norm is upper bounded by $3\mathop{\mathrm{tr}}(O_{0}^{2})$ [10]. Consequently, for all unknown states $\rho$ and observables that permit efficient stabilizer state decomposition, the computational complexity of calculating $\mathop{\mathrm{tr}}(AB)$ achieves exponential speedup, expressed as
$\mathcal{N}\leq O\left(\frac{\log\frac{L}{\sigma}}{\epsilon^{2}}\right)\times 3\,\mathop{\mathrm{tr}}(O_{0}^{2})\times\text{poly}(n),$
where $\text{poly}(n)$ denotes the maximum post-processing complexity in each experiment.
Definition 1 (Approximately DDB-Average State).
A state $\rho$ is called approximately DDB-average if it satisfies the following condition:
$\left|\text{tr}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d},$ (15)
for all $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$.
Lemma 1.
For any bounded-norm observable $O$ and approximately DDB-average state $\rho$, the variance of shadow tomography via random DDB measurements satisfies
$\lVert O_{0}\rVert_{\rho}^{2}\leq O(\text{poly}(\log d))\times\text{tr}(O_{0}^{2}).$
Consequently, the computational complexity of estimating $\mathop{\mathrm{tr}}(\rho O)$ via random DDB measurements is $O(\text{poly}(\log d))$.
Details and proofs are provided in Appendix B.3. We introduce a special state $\rho_{A}$, defined as
$\rho_{A}=\frac{I}{d}+\sum_{j\neq k}\rho_{jk}|j\rangle\langle k|,$ (16)
where $\rho_{jk}=\rho^{*}_{kj}$ and $0\leq|\rho_{jk}|<\frac{O(\text{poly}(\log d))}{d}$. This state is approximately DDB-average, allowing shadow estimation of $\mathop{\mathrm{tr}}(\rho_{A}O)$ with any desired accuracy while maintaining polynomial computational complexity for bounded norm observable $O$. However, $\rho_{A}$ involves $O(d^{2})$ variables, which imposes computational challenges for the same task with only classical computing.
Conclusion and discussion—In this work, we introduce a novel shadow tomography method applicable to arbitrary $d$-level systems as a quantum algorithm for computing $\mathop{\mathrm{tr}}(AB)$, where $A$ is a trace-1 Hermitian density matrix and $B$ is a bounded-norm observable. This approach has potential applications in high-dimensional quantum science, AI, and optimization. For $d$-dimensional matrices $A$ and $B$, classical computational complexity is $O(d^{2})$. By employing shadow tomography with random DDB measurements, we reduce the worst-case complexity to $O(d)$ and achieve $O(\text{poly}(\log d))$ complexity for matrices $A$ that satisfy $\lVert B-\mathop{\mathrm{tr}}(B)B/d\rVert_{A}^{2}=O(\text{poly}(\log d))$. While matrix $A$ may have more restricted applicability compared to random Clifford measurements, our method remains efficient for general bounded-norm matrices $B$ and approximately DDB-average $A$. Additionally, the memory requirement for data storage is reduced to a polynomial amount in $\log d$ classical bits.
There are several promising directions for future research. First, expanding the class of matrices $A$ with polynomial sampling complexity is essential for enhancing the practical applicability. This expansion could involve biased sampling, additional constraints on observables, or exploring alternative snapshot sets with at most two nonzero amplitudes. Notably, the DDB and MUB snapshots include $O(2^{2n})$ elements for $d=2^{n}$, where both worst-case and average variance exhibit similar behavior. Expanding the MUB snapshot set to $O(2^{n^{2}})$ stabilizer states could further reduce worst-case variance to $3\mathop{\mathrm{tr}}(O^{2}_{0})$.
Second, while our work primarily addresses the efficient computation of $\mathop{\mathrm{tr}}(AB)$, extending these techniques to nonlinear properties, such as purity and entropy [18], could broaden the applications.
Third, many current quantum algorithms focus on achieving exponential speedups in Stage One (Fig. 1) using fixed quantum circuits. Incorporating measurements or randomized quantum circuits could unlock new potential for achieving quantum advantage in this phase.
Finally, improving the robustness and accuracy of the scheme in noisy environments is crucial [19, 20, 21]. Since our primary goal is to estimate expectation values rather than fully reconstruct quantum states, this approach requires less extensive error correction than typical quantum computations.
Acknowledgements— The work of Yu Wang received support from the National Natural Science Foundation of China through Grants No. 62001260 and No. 42330707, as well as from the Beijing Natural Science Foundation under Grant No. Z220002.
Appendix A Calculation of reconstruction channel
We may express $\rho$ with the following form
$\rho=\sum_{j,k=0}^{d-1}\rho_{jk}|j\rangle\langle k|.$ (17)
When $d$ is even, there are $12d-$ 1 DDBs, with the computational basis states $\{|0\rangle,|1\rangle,\cdots,|d-1\rangle\}$ sampled twice. When $d$ is odd, there are $12d-$ 1 DDBs but the computational states $\{|t\rangle:t=0,\cdots,d-1\}$ appear twice. Thus for general dimension $d$, the quantum channel for randomly sampling DDBs is given by:
$2d\times\mathcal{M}(\rho)$ $=2\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})P_{k}+\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\rho P_{jk}^{\pm})\cdot P_{jk}^{\pm}+\text{tr}(\rho Q_{jk}^{\pm})\cdot Q_{jk}^{\pm}\right]$ (18)
$=2\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\sum_{0\leq j<k\leq d-1}[(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)+\rho_{jk}|j\rangle\langle k|+\rho_{kj}|k\rangle\langle j|]$ (19)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\sum_{0\leq j<k\leq d-1}(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)$ (20)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}[\sum_{j=0}^{d-1}\sum_{k=0}^{d-1}(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)-\sum_{k=0}^{d-1}(\rho_{kk}+\rho_{kk})(|k\rangle\langle k|+|k\rangle\langle k|)]$ (21)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}[\sum_{j=0}^{d-1}(d\rho_{jj}|j\rangle\langle j|+\rho_{jj}I+\mathop{\mathrm{tr}}(\rho)|j\rangle\langle j|+\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|)-4\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|]$ (22)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}(2d\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+2\mbox{tr}(\rho)I-4\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|)$ (23)
$=\rho+\mbox{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|.$ (24)
From Eq.(18) to Eq.(19), we use
$\begin{cases}\text{tr}(\rho P_{jk}^{+})=(\rho_{jj}+\rho_{kk}+\rho_{jk}+\rho_{kj})/2,\\ \text{tr}(\rho P_{jk}^{-})=(\rho_{jj}+\rho_{kk}-\rho_{jk}-\rho_{kj})/{2},\\ \text{tr}(\rho Q_{jk}^{+})=(\rho_{jj}+\rho_{kk}+i\rho_{jk}-i\rho_{kj})/{2},\\ \text{tr}(\rho Q_{jk}^{-})=(\rho_{jj}+\rho_{kk}-i\rho_{jk}+i\rho_{kj})/{2},\end{cases}$ (25)
and
$\begin{cases}P_{jk}^{+}=(|j\rangle\langle j|+|k\rangle\langle k|+|j\rangle\langle k|+|k\rangle\langle j|)/2,\\ P_{jk}^{-}=(|j\rangle\langle j|+|k\rangle\langle k|-|j\rangle\langle k|-|k\rangle\langle j|)/2,\\ Q_{jk}^{+}=(|j\rangle\langle j|+|k\rangle\langle k|-i|j\rangle\langle k|+i|k\rangle\langle j|)/2,\\ Q_{jk}^{-}=(|j\rangle\langle j|+|k\rangle\langle k|+i|j\rangle\langle k|-i|k\rangle\langle j|)/2.\end{cases}$ (26)
Denote the following symbols.
$\begin{cases}a_{jk}=\rho_{jj}+\rho_{kk},\\ b_{jk}^{\pm}=\rho_{jk}\pm\rho_{kj},\\ A_{jk}=|j\rangle\langle j|+|k\rangle\langle k|\\ B_{jk}^{\pm}=|j\rangle\langle k|\pm|k\rangle\langle j|.\\ \end{cases}$ (27)
Thus we have
$\begin{cases}\mathop{\mathrm{tr}}(\rho P^{+}_{jk})P^{+}_{jk}=\frac{1}{4}(a_{jk}+b_{jk}^{+})(A_{jk}+B_{jk}^{+})\\ \mathop{\mathrm{tr}}(\rho P^{-}_{jk})P^{-}_{jk}=\frac{1}{4}(a_{jk}-b_{jk}^{+})(A_{jk}-B_{jk}^{+})\\ \mathop{\mathrm{tr}}(\rho Q^{+}_{jk})Q^{+}_{jk}=\frac{1}{4}(a_{jk}+ib_{jk}^{-})(A_{jk}-iB_{jk}^{-})\\ \mathop{\mathrm{tr}}(\rho Q^{-}_{jk})Q^{-}_{jk}=\frac{1}{4}(a_{jk}-ib_{jk}^{-})(A_{jk}+iB_{jk}^{-}).\\ \end{cases}$ (28)
It is easy to verify that the summation above is equal to $a_{jk}A_{jk}+(b_{jk}^{+}B^{+}_{jk}+b_{jk}^{-}B^{-}_{jk})/2=(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)+\rho_{jk}|j\rangle\langle k|+\rho_{kj}|k\rangle\langle j|$.
Thus the quantum channel is given by
$\mathcal{M}(\rho)=\frac{1}{2d}[\rho+\mbox{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|].$ (29)
As a comparison, when we use the uniform sampling of Clifford measurements or MUB measurements, the quantum channel is given by $\mathcal{M}(\rho)=\frac{1}{d+1}(\rho+\mathop{\mathrm{tr}}(\rho)I)$, where $d=2^{n}$.
For the channel corresponding to uniform sampling from Cliffords and MUBs, the mapping of the matrix elements of a density matrix $\rho$ is as follows:
  • The off-diagonal elements $\rho_{jk}$ (where $j\neq k$) are mapped to $\rho_{jk}/(d+1)$.
  • The diagonal elements $\rho_{jj}$ are mapped to $\frac{\rho_{jj}+\mathrm{tr}(\rho)}{d+1}$.
In contrast, for the channel corresponding to uniform sampling from DDBs, the mapping is:
  • The off-diagonal elements $\rho_{jk}$ (where $j\neq k$) are mapped to $\rho_{jk}/(2d)$.
  • The diagonal elements $\rho_{jj}$ are mapped to $\frac{d\times\rho_{jj}+\mathrm{tr}(\rho)}{2d}$.
The inverse reconstruction channel of uniform sampling DDBs is given by
$\mathcal{M}^{-1}(\rho)$ $=2d\Bigg{[}\rho-\frac{d-1}{d}\sum_{k=0}^{d-1}\mbox{tr}(\rho P_{k})P_{k}\Bigg{]}-\frac{\mbox{tr}(\rho)}{d}I.$ (30)
It also includes the linear combination of $\rho$ and $I$, with additional corrections to the diagonal terms.
Appendix B Performance guarantee
The predicted observable is $O$. Denote its traceless part as $O_{0}=O-\mbox{tr}(O)I/d$. If we uniformly sample the DDBs for state $\sigma$ as introduced above, the sampling complexity is linearly dependent on the variance in Eq. (4).
We have
$\mathcal{M}^{-1}(O_{0})=2d\left[O_{0}-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(O_{0}P_{k})P_{k}\right].$ (31)
Denote $\mathcal{M}^{-1}(O_{0})=2d\times o$, where $o=O_{0}-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(O_{0}P_{k})P_{k}$.
By the definition of $o$, we can calculate the relationship between its matrix elements and those of the matrix $O_{0}$:
$&\text{tr}(oP_{k})=\frac{1}{d}\text{tr}(O_{0}P_{k})\\ &\text{tr}(o|j\rangle\langle k|)=\text{tr}(O_{0}|j\rangle\langle k|).\\ \end{aligned}\right.$ (32)
This means that the diagonal elements of the operator $o$ correspond to the reciprocal of the diagonal elements of operator $O_{0}$ with a denominator of $d$. The operators $o$ and $O_{0}$ share the same non-diagonal elements. The variance for unknown state $\sigma$ is then expressed as follows:
$\lVert O_{0}\rVert_{\sigma}^{2}$ $=\sum_{k=0}^{d-1}\frac{2\text{tr}(\sigma P_{k})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P_{k})+\sum_{0\leq j<k\leq d-1}[\frac{\text{tr}(\sigma P^{+}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P^{+}_{jk})+\frac{\text{tr}(\sigma Q^{+}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})Q^{+}_{jk})]$ (33)
$+\frac{\text{tr}(\sigma P^{-}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P^{-}_{jk})+\frac{\text{tr}(\sigma Q^{-}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})Q^{-}_{jk})]$
$=4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\sigma P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\sigma Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right].$
B.1 Upper bound for the worst case
We know that $\mbox{tr}(\sigma P_{k}),\mbox{tr}(\sigma P_{jk}^{\pm}),\mbox{tr}(\sigma Q_{jk}^{\pm})\leq 1$. To estimate the upper bound of $\lVert O_{0}\rVert_{\sigma}^{2}$, we can roughly conclude
$\lVert O_{0}\rVert_{\sigma}^{2}\leq 4d\sum_{k=0}^{d-1}\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}^{2}(oQ_{jk}^{\pm})\right]).$ (34)
Denote the matrix elements of $o$ as $o_{jk}$, where $j,k\in\{0,\cdots,d-1\}$. We rewrite Eq. (34) as $\lVert O_{0}\rVert_{\sigma}^{2}\leq 2d\times T$, where
$T\doteq$ $\sum_{k=0}^{d-1}2\text{tr}^{2}(oP_{k})+\sum_{0\leq j<k\leq d-1}\left[\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}^{2}(oQ_{jk}^{\pm})\right]$
$=$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\frac{1}{4}\sum_{0\leq j<k\leq d-1}\big{[}(o_{jj}+o_{kk}+o_{jk}+o_{kj})^{2}+(o_{jj}+o_{kk}-o_{jk}-o_{kj})^{2}$
$+(o_{jj}+o_{kk}+io_{jk}-io_{kj})^{2}+(o_{jj}+o_{kk}-io_{jk}+io_{kj})^{2}\big{]}$
$=$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}\big{[}(o_{jj}+o_{kk})^{2}+2o_{jk}o_{kj}\big{]}$
$\leq$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}\big{[}2(o_{jj}^{2}+o_{kk}^{2})+2o_{jk}o_{kj}\big{]}$
$=$ $2d\sum_{k=0}^{d-1}o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}$
$=$ $\frac{2d}{d^{2}}\sum_{k=0}^{d-1}[d\times o_{kk}]^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}.$
When $d\geq 2$, we have $\frac{2d}{d^{2}}\leq 1$. So, we can deduce the upper bound of $T$.
$T\leq\sum_{k=0}^{d-1}[d\times o_{kk}]^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}=\text{tr}(O_{0}^{2}).$
Here we use the relation in Eq. (32).
Thus the upper bound of $\lVert O_{0}\rVert_{\sigma}^{2}$ can be deduced. For each unknown state $\sigma$, the expectation values of $\text{tr}(\sigma P_{k})$, $\text{tr}(\sigma P_{jk}^{\pm})$, and $\text{tr}(\sigma Q_{jk}^{\pm})$ are no bigger than 1. Then we have
$\lVert O_{0}\rVert_{\text{ shadow}}^{2}=\max_{\sigma:\text{ state }}\lVert O_{0}\rVert_{\sigma}^{2}\leq 2dT\leq 2d\times\text{tr}(O_{0}^{2}).$ (35)
Consider the variance in Eq. (33), the first part $4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})$ tends to zero as $d$ increases. As we have $\text{tr}^{2}(oP_{k})=\frac{\mathop{\mathrm{tr}}^{2}(O_{0}|k\rangle\langle k|)}{d^{2}}\leq\mathop{\mathrm{tr}}(O^{2}_{0})/d^{2}$ for all $k$. Thus
$4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})\leq\frac{4\mathop{\mathrm{tr}}(O^{2}_{0})}{d}.$ (36)
Then for bounded-norm observable, the efficiency of sampling complexity is just related to the following part:
$V_{\mbox{diag}}=2d\times\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\sigma P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\sigma Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right].$ (37)
One worst case could happen when the unknown state and the observable are the same as one of the nontrivial DDB states in Eq.(9). For example, $\sigma=P_{01}^{+}$ and $O=P_{01}^{+}$. Then $O_{0}=P_{01}^{+}-I/d$, $o=\frac{d-2}{2d^{2}}(|0\rangle\langle 0|+|1\rangle\langle 1|)+\frac{|0\rangle\langle 1|+|1\rangle\langle 0|}{2}$. Thus $\mathop{\mathrm{tr}}(oP_{jk})>1/2$ and $V_{\mbox{diag}}>d$. Then in this case, the sampling complexity is linear dependent with $d$.
B.2 Average performance analysis
If we sample $|\phi\rangle=U|0\rangle$ with Haar measure, the average state will be
$\int_{U(2^{n})}U|0\rangle\langle 0|U^{{\dagger}}\mathrm{d}\mu(U)=I/2^{n}\;.$ (38)
Thus the variance of state $\sigma=I/d$ exhibits the average performance when the output $\rho$ is randomly and uniformly generated.
When the unknown state is $\sigma=I/d$, we have
$\mathop{\mathrm{tr}}(\sigma P_{k})=\mathop{\mathrm{tr}}(\sigma P^{\pm}_{jk})=\mathop{\mathrm{tr}}(\sigma Q^{\pm}_{jk})=1/d.$ (39)
Take Eq. (39) into Eq. (33), we have
$\lVert O_{0}\rVert_{I/d}^{2}\leq\frac{1}{d}2dT\leq 2\text{tr}(O_{0}^{2}).$ (40)
Thus the average performance is efficient for bounded norm observables. This is constant times of the average performance when we uniformly sample MUB measurements, which is given by $(1+1/2^{n})\text{tr}(O_{0}^{2})$, where $d=2^{n}$. However, the informationally complete $d+1$ MUBs have only been found in prim power dimensions.
B.3 Approximate average case
If the state $\rho$ is approximately DDB-average, then its deviation from the completely mixed state $I/d$ is small. Specifically, we have
$\left|\text{tr}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d}$
for all snapshots $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$. This implies that $\text{tr}(\rho|\phi\rangle\langle\phi|)\leq\frac{O(\text{poly}(\log d))+1}{d}$. Substituting this into Eq. (33), we can deduce
$\lVert O_{0}\rVert_{\rho}^{2}$ $=4d\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})\cdot\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\rho P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\rho Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right]$ (41)
$\leq O(\text{poly}(\log d)+1)\times 2T$
$\leq O(\text{poly}(\log d))\times\text{tr}(O^{2}_{0}).$
Denote $\epsilon_{1}=\frac{O(\text{poly}(\log d))}{d}$. It is an interesting question to characterize the proportion of randomly chosen states for various levels of deviation $\epsilon_{1}$, such as $\frac{\log d}{d}$, $\frac{\log^{2}d}{d}$, $\frac{\log^{3}d}{d}$, and so forth.