### v4

# Quantum Advantage via Efficient Post-processing on Qudit Shadow Tomography

Abstract
The computation of $\operatorname{tr}(AB)$ is essential in quantum science and artificial intelligence, yet classical methods for $d$-dimensional matrices $A$ and $B$ require $O(d^{2})$ complexity, which becomes infeasible for exponentially large systems. We introduce a quantum approach based on qudit shadow tomography that reduces both computational and storage complexities to $O(\text{poly}(\log d))$ in specific cases. This method is applicable to quantum density matrices $A$ and Hermitian matrices $B$ with given $\operatorname{tr}(B)$ and $\operatorname{tr}(B^{2})$ bounded by a constant (referred to as BN-observables). We prove that this method guarantees at least a quadratic speedup for any quantum state $\rho$ and BN-observable $O$ in the worst case, and an exponential speedup in the approximately average case. For any $n$-qubit stabilizer state $\rho$ and arbitrary BN-observable $O$, we show that $\operatorname{tr}(\rho O)$ can be efficiently estimated with $\text{poly}(n)$ computations. Moreover, our approach significantly reduces the post-processing complexity in shadow tomography using random Clifford measurements, and it is applicable to arbitrary dimensions $d$. These advances open new avenues for efficient high-dimensional data analysis and modeling.
Introduction—Quantum advantage represents a transformative leap in computational power, breaking classical limits across various fields. From the exponential query complexity reduction in the Deutsch-Jozsa algorithm [1] to Shor’s algorithm for breaking RSA encryption [2] and the HHL algorithm’s ability to solve linear equations exponentially faster [3], quantum algorithms continue to reshape what is computationally feasible. Recent advancements, such as random circuit sampling on NISQ devices [4, 5], underscore the potential for groundbreaking applications. With increasing demands for high-dimensional data processing, particularly in artificial intelligence, advancing quantum advantage holds transformative potential.
Efficiently computing $\operatorname{tr}(AB)$ is crucial in quantum science, machine learning, optimization, and data analysis, particularly for tasks involving matrix operations, quantum measurements, and kernel methods [6, 7]. However, as the dimension $d$ increases exponentially (e.g., $d=2^{n}$), both storing matrices and computing traces become infeasible. In contrast, an $n$-qubit quantum state can represent a $2^{n}$-dimensional density matrix, significantly reducing storage needs in quantum simulations [8].
We consider Hermitian, positive semi-definite matrices $A$ that represent quantum density matrices (trace 1), and Hermitian matrices $B$ with specified $\operatorname{tr}(B)$ and bounded $\operatorname{tr}(B^{2})$ (BN-observables). Classically, computing $\operatorname{tr}(AB)$ requires $O(d^{2})$ complexity, which is impractical for exponentially high-dimensional systems. Utilizing quantum advantage, we achieve exponential speedup for approximately average cases in both computation and storage, allowing polynomial-time evaluation of $\operatorname{tr}(AB)$ with arbitrary precision, and guaranteeing at least quadratic speedup in the worst case. For any $n$-qubit stabilizer state and BN-observable $O$, we show that $\operatorname{tr}(\rho O)$ can be computed in polynomial time. This method is particularly effective for high-dimensional problems where classical approaches fail, expanding quantum computing’s reach to tackle complex tasks, as shown in Fig. 1.
Figure 1: Advantage seeking: In the absence of noise, the output of quantum computation is an $n$-qubit pure state $|x\rangle$, representing a $2^{n}$-dimensional vector. This state can serve as an intermediate step in quantum machine learning algorithms, such as solving $A\vec{x}=\vec{b}$ [3], and is commonly used as a sub-procedure in various quantum machine learning algorithms [9]. We assume that generating $|x\rangle$ through quantum computation is not exponentially slower than classical counterparts, though many quantum algorithms seek exponential speedup at this stage. All quantum states are represented by a density matrix $\rho$. Although $\rho$ has a limited lifetime, it can be stored as polynomial $n$ classical bits through well-designed random measurements. The main challenge is determining whether we can efficiently compute $\text{tr}(\rho O_{k})$ from these $n$ bits with polynomial-time complexity and arbitrary precision. Additionally, we explore whether this approach can be extended to a broader set of density matrices $\rho$ for BN-observables $O$.
We build on the shadow tomography framework [10], a key method for efficiently estimating properties $\{\operatorname{tr}(\rho O_{k})\}$ of an unknown quantum state $\rho$. Shadow tomography provides a significant quantum advantage in experimental learning by reducing sample complexity exponentially compared to classical methods [11]. Calculating $\operatorname{tr}(\rho O)$ for a BN-observable $O$ is essential for tasks like fidelity estimation and entanglement detection. However, existing methods based on random Clifford and MUB measurements [12] face challenges. Even with polynomial sample complexity, the post-processing complexity for general observables without efficient expression can still be exponential. This issue is common in current $n$-qubit shadow tomography frameworks when dealing with general observables. Moreover, for arbitrary $d$-level systems, processing Clifford measurements and finding $d+1$ MUBs remain unresolved problems [13]. On optical platforms, projecting onto random states instead of using full Clifford or MUB circuits for $n$-qubit systems results in a time complexity of $O(2^{n}n^{3})$ [14]. These issues highlight the need for alternative approaches to enhance the efficiency and applicability of shadow tomography across different quantum systems.
In this work, we introduce DDB-ST (DDB-based Shadow Tomography), a new shadow tomography scheme based on random projective measurements onto $2d$ Dense Dual Bases (DDBs). Unlike Clifford-ST (Clifford-based Shadow Tomography), which can involve exponential post-processing per experiment for the worst case, DDB-ST requires only constant post-processing per measurement, enabling efficient polynomial-time computation as long as the sample complexity remains polynomial.
We show that this sample complexity of DDB-ST is independent of the system dimension $d$ on average, and scales as $O(d)$ in the worst case, offering a consistent quadratic speedup for computing $\operatorname{tr}(\rho O)$. In $n$-qubit systems, applying DDB-ST to stabilizer states enables polynomial-time computation of the expectation value for any BN-observable. DDB measurements are particularly well-suited for optical systems, where projected states have at most two non-zero amplitudes, making them both experimentally feasible and efficient.
Preliminaries and framework—Shadow tomography [10] is to predict an unknown quantum state’s properties efficiently. The process involves randomly performing operations from an informationally complete unitary ensemble $\mathcal{U}=\{U_{k}\}$. In each experiment, the state $\rho$ is evolved using $U_{k}$ with probability $p_{k}$, measured in the computational basis, and then the result $j$ is recorded. It corresponds to a quantum process
$\mathcal{M}(\rho)=\sum_{k}p_{k}\mathop{\mathrm{tr}}(U_{k}\rho U_{k}^{{\dagger}}|j\rangle\langle j|)U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}.$ (1)
Each experiment yields a single estimate of the unknown observable $O$ as
$\mathop{\mathrm{tr}}(O\mathcal{M}^{-1}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k})).$ (2)
Let $\langle o_{k}\rangle$ as the final estimation of $\mathop{\mathrm{tr}}(\rho O_{k})$. Since the quantum state collapses upon measurement, multiple copies of the unknown state should be prepared. Given exponentially large number of observables $\{O_{k}\}_{k=1}^{L}$, to achieve accuracy $\epsilon$ and confidence level $1-\sigma$, we can ensure
$P_{r}(|\langle o_{k}\rangle-\mathop{\mathrm{tr}}(\rho O_{k})|<\epsilon)\geq 1-\sigma$ (3)
with the following sample complexity:
$N=O\left(\frac{\log\frac{L}{\sigma}}{\epsilon^{2}}\right)\max_{1\leq i\leq L}\left\|O_{i}-\frac{\text{tr}(O_{i})}{d}I\right\|_{\mbox{shadow}}^{2},$ (4)
where $\|\cdot\|_{\mbox{shadow}}^{2}=\max_{\sigma:\text{ state }}\|\cdot\|_{\sigma}^{2}$ denotes the shadow norm associated with $\mathcal{U}$ and observable $O$. It represents the worst-case sample complexity, while $\|\cdot\|_{\sigma}^{2}$ is related to the sample complexity for state $\sigma$. Denote $O_{0}=O-\mathop{\mathrm{tr}}(O)I/d$, $\lVert O_{0}\rVert_{\sigma}^{2}$ is
$\mathbb{E}_{U\sim\mathcal{U}}\sum_{b\in\{0,1\}^{n}}\langle b|U\sigma U^{\dagger}|b\rangle\cdot\langle b|U\mathcal{M}^{-1}(O_{0})U^{\dagger}|b\rangle^{2}.$ (5)
If the ensemble $\mathcal{U}$ consists of all $n$-qubit Clifford circuits, the reconstruction channel $\mathcal{M}^{-1}$ is given by:
$\mathcal{M}^{-1}(X)=(2^{n}+1)X-\operatorname{tr}(X)I.$
The corresponding variance satisfies $\|\cdot\|_{\text{shadow}}\leq 3\operatorname{tr}^{2}(O_{0})$. Hence, polynomial sampling of the unknown state is sufficient for predicting BN-observables by using Clifford-ST.
Essentially, shadow tomography can be viewed as a quantum algorithm to approximate $\mathop{\mathrm{tr}}(\rho O)$ with arbitrary accuracy. For BN-observable $O$ and any density matrix $\rho$, the estimation using Clifford-ST is given by:
$\mathop{\mathrm{tr}}(\rho O)\approx\sum_{k=1}^{\text{poly}(n)}\frac{(2^{n}+1)\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)-2^{n}\mathop{\mathrm{tr}}(O)}{\text{poly}(n)}.$ (6)
Here, $\text{poly}(n)$ corresponds to the sample complexity in Eq. (4). Statistical methods like the median of means are used to refine data reconstruction with partial data, but they do not reduce the overall sample complexity.
The overall computational complexity remains polynomial if each calculation of $\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)$ is constant or polynomial rather than exponential. The snapshots, or projected states, form the set
$\mathcal{S}=\{U_{k}^{{\dagger}}|j\rangle:U_{k}\in\mathcal{U},j=0,\cdots,d-1\}.$ (7)
However, the snapshots generated by random Clifford measurements are all stabilizer states. If the observable $O$ lacks an efficient representation in terms of states in $\mathcal{S}$, the calculation of $\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)$ in a single experiment can become exponential, posing challenges similar to those in the classical computation of $\mathop{\mathrm{tr}}(AB)$. This reflects a common issue related to the exponential post-processing complexity in $n$-qubit shadow tomography.
To address this, we modify the collection of snapshots to ensure that the computational complexity of each calculation $\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}O)$ remains constant for any $O$.
Snapshots with dense dual basis states— We define the following states:
$\begin{cases}|\phi_{jk}^{\pm}\rangle=\frac{1}{\sqrt{2}}(|j\rangle\pm|k\rangle),\\ |\psi_{jk}^{\pm}\rangle=\frac{1}{\sqrt{2}}(|j\rangle\pm i|k\rangle).\end{cases}$ (8)
The new collection of snapshots comprises a total of $2d^{2}-d$ elements:
$\mathcal{S}_{\text{DDB}}=\{P_{t}=|t\rangle\langle t|,~{}~{}t=0,\ldots,d-1;\\ P_{jk}^{\pm}=|\phi_{jk}^{\pm}\rangle\langle\phi_{jk}^{\pm}|,Q_{jk}^{\pm}=|\psi_{jk}^{\pm}\rangle\langle\psi_{jk}^{\pm}|;~{}~{}0\leq j<k\leq d-1\}.$ (9)
These rank-1 projections are informationally complete and thus sufficient for shadow tomography.
An efficient algorithm with $O(\log d)$ time complexity has been developed to construct a unitary ensemble $\mathcal{U}=\{U_{j}\}_{j=1}^{f(d)}$ with the minimal number of elements required to span all states in Eq. (9) [15]. Each orthonormal basis $\{U_{j}|k\rangle:k=0,\dots,d-1\}$ is referred to as a Dense Dual Basis (DDB).
For even $d$, $f(d)=2d-1$. The first DDB corresponds to the computational basis, while the remaining bases are constructed in dual pairs, denoted as $\{|\phi_{jk}^{\pm}\rangle:(j,k)\in\mathbb{T}\}$ and $\{|\psi_{jk}^{\pm}\rangle:(j,k)\in\mathbb{T}\}$, where $\mathbb{T}$ is a partition of $\{0,1,\dots,d-1\}$ into distinct pairs with no repeated elements. For example, when $d=4$, one possible partition is $\mathbb{T}_{1}=\{(0,1),(2,3)\}$. A minimum of $d-1$ such partitions is needed to cover all pairs $(j,k)$ where $10\leq j<k\leq d-$ 1.
For odd $d$, $f(d)=2d$. In this case, the DDBs are similarly paired as $\{|\phi_{jk}^{\pm}\rangle,|l_{\mathbb{T}}\rangle:(j,k)\in\mathbb{T}\}$ and $\{|\psi_{jk}^{\pm}\rangle,|l_{\mathbb{T}}\rangle:(j,k)\in\mathbb{T}\}$, where $l_{\mathbb{T}}$ represents the single element in $\{0,1,\dots,d-1\}$ not included in $\mathbb{T}$. For example, with $d=5$ and $\mathbb{T}_{1}=\{(0,1),(2,3)\}$, the value of $l_{\mathbb{T}}$ is 4. A minimum of $d-1$ such partitions is also sufficient.
In $n$-qubit systems, each DDB circuit can be implemented using one Hadamard gate (or followed by phase gate $S$) and a permutation gate, with a decomposition into $O(n^{4})$ elementary gates.
Compared with random Clifford measurements for shadow tomography, random DDB measurements offer several benefits. First, they are applicable across arbitrary dimensions. Second, in optical systems with random snapshot projections, the computational cost is significantly reduced. For instance, calculating a random stabilizer state requires $O(2^{n}n^{3})$ operations for dimension $d=2^{n}$ [14], whereas DDB states contain at most two nonzero amplitudes, greatly reducing computational demand. Third, the biased sampling process is more efficient; while the total number of stabilizer states scales as $O(2^{n^{2}})$, DDB states number only $O(2^{2n})$. Additionally, DDB-ST can also yield an explicit expression for the reconstruction channel.
Theorem 1 (Reconstruction channel).
For odd dimensions $d$, each unitary operation in the DDB ensemble $\{U_{j}\}_{j=1}^{2d}$ is sampled with probability $1/2d$. For even dimensions $d$, the ensemble $\{U_{j}\}_{j=1}^{2d-1}$ is sampled, with a probability of $2/2d$ assigned to $I$ (for computational basis states) and $1/2d$ to all other elements. The resulting quantum channel is
$\mathcal{M}(\rho)=\frac{1}{2d}\left[\rho+\text{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|\right].$ (10)
The corresponding reconstruction channel is
$\mathcal{M}^{-1}(\rho)=2d\left[\rho-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})P_{k}\right]-\frac{\text{tr}(\rho)}{d}I.$ (11)
The proof is detailed in Appendix A. In each single experiment, the calculation of $\mathop{\mathrm{tr}}\left[\mathcal{M}^{-1}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k})O\right]$ during post-processing becomes a constant. This efficiency arises because the term $U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}-\frac{d-1}{d}\sum_{k=0}^{d-1}\mathop{\mathrm{tr}}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}P_{k})P_{k}$ contains at most four nonzero elements, and $\mathop{\mathrm{tr}}(O)$ is known.
Theorem 2 (Performance guarantee).
In a $d$-dimensional Hilbert space, if we use shadow tomography with DDB measurements for predicting the expectation value of a given observable $O$, worst-case variance for a quantum state $\sigma$ is bounded by
$\lVert O_{0}\rVert_{\text{shadow}}^{2}=\max_{\sigma:\text{ state}}\lVert O_{0}\rVert_{\sigma}^{2}\leq 2d\,\text{tr}(O_{0}^{2}),$ (12)
where $O_{0}=O-\mathop{\mathrm{tr}}(O)/d$. If the unknown state is sampled randomly according to the Haar measure, the average variance is bounded by
$\lVert O_{0}\rVert_{I/d}^{2}\leq 2\text{tr}(O_{0}^{2})$ (13)
The detailed proof is provided in Appendix B. For an arbitrary unknown quantum state $\rho$ (represented as a Hermitian, positive semi-definite matrix $A$ with $\mathop{\mathrm{tr}}(A)=1$), the computational complexity of estimating $\mathop{\mathrm{tr}}(\rho O)$ using DDB-ST is:
$\mathcal{N}_{\text{DDB}}\leq O\left(\frac{\log\frac{1}{\sigma}}{\epsilon^{2}}\right)\times\lVert O_{0}\rVert_{\rho}^{2}\times\text{constant},$ (14)
where $\epsilon$ is the error tolerance, and $1-\sigma$ is the confidence level.
Thus with DDB-ST, we achieve an exponential speedup in computing $\mathop{\mathrm{tr}}(AB)$ (or $\mathop{\mathrm{tr}}(\rho O)$) for all states $\rho$ and BN-observables $O$ that satisfy $\lVert O_{0}\rVert_{\rho}^{2}=O(\text{poly}(\log d))$. In the worst-case scenario, at least a quadratic speedup is guaranteed due to the bound $\lVert O_{0}\rVert_{\text{shadow}}^{2}\leq 2d\mathop{\mathrm{tr}}(O_{0}^{2})$. On average, polynomial complexity can be achieved when $\rho$ is randomly chosen.
In contrast, the shadow norm by Clifford-ST is upper bounded by $3\mathop{\mathrm{tr}}(O_{0}^{2})$ [10]. Consequently, for all unknown states $\rho$ and specific BN-observables $O$ that permit efficient stabilizer state decomposition, the computational complexity of calculating $\mathop{\mathrm{tr}}(AB)$ could achieve exponential speedup, expressed as
$\mathcal{N}_{\text{Clifford}}\leq O\left(\frac{\log\frac{1}{\sigma}}{\epsilon^{2}}\right)\times 3\,\mathop{\mathrm{tr}}(O_{0}^{2})\times\text{poly}(n),$ (15)
where $\text{poly}(n)$ denotes the maximum post-processing complexity in the experiment.
Definition 1 (Approximately DDB-Average State).
A state $\rho$ is called approximately DDB-average if it satisfies the following condition:
$\left|\text{tr}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d},$ (16)
for all $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$.
We conducted numerical experiments to analyze the proportion of approximately DDB-average states, as shown in Fig. (2). For dimensions $d=2^{2},\ldots,2^{8}$, 1000 rank-1 density matrices $\rho$ were randomly generated according to the Haar measure. For each $\rho$, we calculated $|\mathop{\mathrm{tr}}(\rho|\phi\rangle\langle\phi|)-1/d|$ for all $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$, recording $\rho$ as approximately DDB-average if the maximum deviation was less than or equal to $s/2^{n}$.
The green bars in Fig. (2) represent the proportion of states with a maximum deviation less than $s/2^{n}$ for $s=4$, the orange bars for $s=2n$, and the blue bars for $s=n^{2}$. The results indicate that as the threshold increases to $s=2n$, the proportion of approximately DDB-average states approaches 100% for most dimensions and reaches exactly 100% when $s=n^{2}$.
Interestingly, in shadow tomography based on uniform sampling of MUBs, the average variance is $(1+1/2^{n})\mathrm{tr}(O_{0}^{2})$, which is approximately half the variance of DDB-ST. For low qubit numbers, setting $s=n$ or $s=2n$ ensures that the randomly generated rank-1 density matrices are approximately MUB-average and approximately DDB-average, respectively. In the case of MUBs, $s=n$ also corresponds to half the value of $s=2n$ in DDB.
Figure 2: Proportion of approximately DDB-average states with $\epsilon=s/2^{n}$ among 1000 randomly sampled quantum states generated according to the Haar measure in $n=2,3,\ldots,8$-qubit systems.
Lemma 1.
For any BN-observable $O$ and an approximately DDB-average state $\rho$, the variance of DDB-ST satisfies:
$\lVert O_{0}\rVert_{\rho}^{2}\leq O(\operatorname{poly}(\log d))\cdot\operatorname{tr}(O_{0}^{2}).$
As a result, the sample and computational complexity for estimating $\operatorname{tr}(\rho O)$ using DDB-ST is $O(\operatorname{poly}(\log d))$.
Details and proofs are provided in Appendix B.3.
Any $n$-qubit stabilizer state [16, 17] is expressed by
$|\Psi\rangle=\frac{1}{\sqrt{2^{r}}}\sum_{k\in A}\omega^{q(k)}|k\rangle,$ (17)
where $\omega^{q(k)}$ is a phase factor taking values in $\{1,-1,i,-i\}$, and
$A=\{Ru+t\mid u\in\mathbb{Z}_{2}^{r}\},$
with $R$ being a nonsingular $N\times r$ matrix, and $t\in\mathbb{Z}_{2}^{N}$ a fixed vector.
Property 1.
For any stabilizer state $|\Psi\rangle$ expressed in Eq. (17), we have
$\max_{|\phi\rangle\in\mathcal{S}_{\text{DDB}}}\text{tr}(|\Psi\rangle\langle\Psi|\phi\rangle\langle\phi|)\leq\frac{1}{2^{r}}.$ (18)
Lemma 2.
Given arbitrary $n$-qubit stabilizer state $|\Psi\rangle$ in Eq. (17), we can efficiently construct a Clifford circuit $T$, composed of elementary gates from the set $\{S,\text{CZ},\text{CX}\}$, such that $|\Phi\rangle=T|\Psi\rangle=|0\rangle^{\otimes(n-r)}\otimes|\Phi_{r}\rangle$. Here, CZ and CX denote the controlled-Z and controlled-X gates, respectively, and $|\Phi_{r}\rangle$ is an $r$-qubit stabilizer state with $2^{r}$ nonzero amplitudes.
Theorem 3.
For any $n$-qubit stabilizer state $|\Psi\rangle$ and any BN-observable $O$, the expectation value $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ can be efficiently estimated using DDB-ST. This estimation requires $O(\operatorname{poly}(n))$ samples and computational resources.
The proofs of Lemma 2 and Theorem 3 are provided in Appendix B.5 and B.6, respectively.
This conclusion does not contradict the Gottesman-Knill theorem. According to the Gottesman-Knill theorem [18], if $O$ is an $n$-qubit Pauli or Clifford operator, the value $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ can be computed efficiently. In particular, for any $n$-qubit Pauli observable $O$, we have $\mathrm{tr}(O^{2})=2^{n}$. However, for a general observable $O$ that cannot be expressed as a concise sum of Clifford operators, the computation becomes exponentially difficult. Lemma 2 applies to all observables $O$ where $\mathrm{tr}(O^{2})$ is bounded by a constant.
In addition to this exponential speedup for estimating $\operatorname{tr}(\rho O)$ for $n$-qubit stabilizer states and arbitrary BN-observables $O$, we consider another scenario where exponential speedup can be achieved by DDB-ST over classical methods.
We introduce a special state $\rho_{A}$, defined as:
$\rho_{A}=\frac{I}{d}+\sum_{j\neq k}\rho_{jk}|j\rangle\langle k|,$
where $\rho_{jk}=\rho^{*}_{kj}$ and $0\leq|\rho_{jk}|<\frac{O(\text{poly}(\log d))}{d}$. This state is approximately DDB-average, allowing efficient shadow estimation of $\operatorname{tr}(\rho_{A}O)$ with arbitrary precision while maintaining polynomial computational complexity for any BN-observable $O$. We can show that: $\left|\operatorname{tr}(\rho_{A}O)-\frac{\operatorname{tr}(O)}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d}\sum_{j\neq k}|O_{jk}|\leq\text{poly}(\log d)\,\sqrt{\operatorname{tr}(O^{2})}.$ However, $\rho_{A}$ and $O$ each involve $O(d^{2})$ variables, making it exponentially hard to perform the same estimation task using purely classical computations.
For any quantum state $\rho$ and BN-observable $O$, the estimator $\operatorname{tr}(\rho O)$ in Eq. (6) can be obtained using Clifford-ST. However, the worst-case scenario involves exponential post-processing complexity due to the need to compute expectation values for stabilizer states $|\Psi\rangle$. While DDB-ST could theoretically be used as a subroutine for this estimation, it does not guarantee polynomial-time computation of $\operatorname{tr}(\rho O)$ with arbitrary precision. The main challenge lies in the fact that, although DDB-ST can estimate $\operatorname{tr}(|\Psi\rangle\langle\Psi|O)$ with precision $\epsilon$ using polynomial samples, the exponential term $12^{n}+$ 1 in Eq. (6) requires controlling the precision of DDB-ST to within $\epsilon\times\text{poly}(n)/2^{n}$. Consequently, this leads to an exponential number of samples and computations when using DDB-ST as a subroutine, resulting in exponential complexity even when combined with Clifford-ST in the worst-case scenario.
Conclusion and discussion— In this work, we introduce a novel shadow tomography method applicable to arbitrary $d$-level systems, providing a quantum algorithm for computing $\operatorname{tr}(AB)$, where $A$ is a trace-1 Hermitian density matrix and $B$ is a BN-observable. This approach has potential applications in high-dimensional quantum science, artificial intelligence, and optimization. For $d$-dimensional matrices $A$ and $B$, classical computation requires $O(d^{2})$ complexity. Using DDB-ST, we reduce the worst-case computational complexity to $O(d)$ and achieve $O(\text{poly}(\log d))$ complexity for matrices $A$ and $B$ that satisfy $\lVert B-\operatorname{tr}(B)B/d\rVert_{A}^{2}=O(\text{poly}(\log d))$. While the sample efficiency of DDB-ST may be more limited compared to Clifford-ST, we demonstrate its effectiveness for approximately DDB-average matrices $A$. Additionally, the post-processing of DDB-ST remains constant, whereas Clifford-ST can become exponentially complex for general $n$-qubit BN-observables in a single experiment. Moreover, we show that the computational complexity can be exponentially reduced for $n$-qubit stabilizer states and general BN-observables $B$. The data storage requirements decreased to a polynomial amount in $\log d$ classical bits correspondingly.
There are several promising avenues for future research. First, expanding the class of matrices $A$ that allow for polynomial sampling complexity is crucial for enhancing practical applicability. This could involve biased sampling, imposing additional constraints on observables, or exploring alternative snapshot sets with at most two nonzero amplitudes. In $n$-qubit systems, DDB measurements and MUB measurements represent two distinct extremes of Clifford measurements. Beyond the computational basis states, DDB states correspond to stabilizer states with $r=2$, while MUB states correspond to stabilizer states with $r=n$. Both DDB and MUB snapshots contain $O(2^{2n})$ elements, exhibiting similar behavior in terms of worst-case and average variance. Maybe the inclusion of more elements with only two nonzero amplitudes could reduce the worst-case sample complexity.
Second, while our work primarily focuses on the efficient computation of $\mathop{\mathrm{tr}}(AB)$, extending these techniques to nonlinear properties, such as purity and entropy [19], could significantly broaden their range of applications.
Third, many current quantum algorithms focus on achieving exponential speedups in Stage One (Fig. 1) using fixed quantum circuits. Incorporating measurements or randomized quantum circuits could unlock new potential for achieving quantum advantage.
Finally, improving the robustness and accuracy of the scheme in noisy environments is crucial [20, 21, 22]. Since our primary goal is to estimate expectation values rather than fully reconstruct quantum states, the task would require less extensive error correction than typical quantum computations.
Acknowledgements— This work received support from the National Natural Science Foundation of China through Grants No. 62001260 and No. 42330707, as well as from the Beijing Natural Science Foundation under Grant No. Z220002.
Appendix A Calculation of reconstruction channel
We may express $\rho$ with the following form
$\rho=\sum_{j,k=0}^{d-1}\rho_{jk}|j\rangle\langle k|.$ (19)
When $d$ is even, there are $12d-$ 1 DDBs, with the computational basis states $\{|0\rangle,|1\rangle,\cdots,|d-1\rangle\}$ sampled twice. When $d$ is odd, there are $12d-$ 1 DDBs but the computational states $\{|t\rangle:t=0,\cdots,d-1\}$ appear twice. Thus for general dimension $d$, the quantum channel for randomly sampling DDBs is given by:
$2d\times\mathcal{M}(\rho)$ $=2\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})P_{k}+\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\rho P_{jk}^{\pm})\cdot P_{jk}^{\pm}+\text{tr}(\rho Q_{jk}^{\pm})\cdot Q_{jk}^{\pm}\right]$ (20)
$=2\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\sum_{0\leq j<k\leq d-1}[(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)+\rho_{jk}|j\rangle\langle k|+\rho_{kj}|k\rangle\langle j|]$ (21)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\sum_{0\leq j<k\leq d-1}(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)$ (22)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}[\sum_{j=0}^{d-1}\sum_{k=0}^{d-1}(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)-\sum_{k=0}^{d-1}(\rho_{kk}+\rho_{kk})(|k\rangle\langle k|+|k\rangle\langle k|)]$ (23)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}[\sum_{j=0}^{d-1}(d\rho_{jj}|j\rangle\langle j|+\rho_{jj}I+\mathop{\mathrm{tr}}(\rho)|j\rangle\langle j|+\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|)-4\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|]$ (24)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}(2d\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+2\mbox{tr}(\rho)I-4\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|)$ (25)
$=\rho+\mbox{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|.$ (26)
From Eq.(20) to Eq.(21), we use
$\begin{cases}\text{tr}(\rho P_{jk}^{+})=(\rho_{jj}+\rho_{kk}+\rho_{jk}+\rho_{kj})/2,\\ \text{tr}(\rho P_{jk}^{-})=(\rho_{jj}+\rho_{kk}-\rho_{jk}-\rho_{kj})/{2},\\ \text{tr}(\rho Q_{jk}^{+})=(\rho_{jj}+\rho_{kk}+i\rho_{jk}-i\rho_{kj})/{2},\\ \text{tr}(\rho Q_{jk}^{-})=(\rho_{jj}+\rho_{kk}-i\rho_{jk}+i\rho_{kj})/{2},\end{cases}$ (27)
and
$\begin{cases}P_{jk}^{+}=(|j\rangle\langle j|+|k\rangle\langle k|+|j\rangle\langle k|+|k\rangle\langle j|)/2,\\ P_{jk}^{-}=(|j\rangle\langle j|+|k\rangle\langle k|-|j\rangle\langle k|-|k\rangle\langle j|)/2,\\ Q_{jk}^{+}=(|j\rangle\langle j|+|k\rangle\langle k|-i|j\rangle\langle k|+i|k\rangle\langle j|)/2,\\ Q_{jk}^{-}=(|j\rangle\langle j|+|k\rangle\langle k|+i|j\rangle\langle k|-i|k\rangle\langle j|)/2.\end{cases}$ (28)
Denote the following symbols.
$\begin{cases}a_{jk}=\rho_{jj}+\rho_{kk},\\ b_{jk}^{\pm}=\rho_{jk}\pm\rho_{kj},\\ A_{jk}=|j\rangle\langle j|+|k\rangle\langle k|\\ B_{jk}^{\pm}=|j\rangle\langle k|\pm|k\rangle\langle j|.\\ \end{cases}$ (29)
Thus we have
$\begin{cases}\mathop{\mathrm{tr}}(\rho P^{+}_{jk})P^{+}_{jk}=\frac{1}{4}(a_{jk}+b_{jk}^{+})(A_{jk}+B_{jk}^{+})\\ \mathop{\mathrm{tr}}(\rho P^{-}_{jk})P^{-}_{jk}=\frac{1}{4}(a_{jk}-b_{jk}^{+})(A_{jk}-B_{jk}^{+})\\ \mathop{\mathrm{tr}}(\rho Q^{+}_{jk})Q^{+}_{jk}=\frac{1}{4}(a_{jk}+ib_{jk}^{-})(A_{jk}-iB_{jk}^{-})\\ \mathop{\mathrm{tr}}(\rho Q^{-}_{jk})Q^{-}_{jk}=\frac{1}{4}(a_{jk}-ib_{jk}^{-})(A_{jk}+iB_{jk}^{-}).\\ \end{cases}$ (30)
It is easy to verify that the summation above is equal to $a_{jk}A_{jk}+(b_{jk}^{+}B^{+}_{jk}+b_{jk}^{-}B^{-}_{jk})/2=(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)+\rho_{jk}|j\rangle\langle k|+\rho_{kj}|k\rangle\langle j|$.
Thus the quantum channel is given by
$\mathcal{M}(\rho)=\frac{1}{2d}[\rho+\mbox{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|].$ (31)
As a comparison, when we use the uniform sampling of Clifford measurements or MUB measurements, the quantum channel is given by $\mathcal{M}(\rho)=\frac{1}{d+1}(\rho+\mathop{\mathrm{tr}}(\rho)I)$, where $d=2^{n}$.
For the channel corresponding to uniform sampling from Cliffords and MUBs, the mapping of the matrix elements of a density matrix $\rho$ is as follows:
  • The off-diagonal elements $\rho_{jk}$ (where $j\neq k$) are mapped to $\rho_{jk}/(d+1)$.
  • The diagonal elements $\rho_{jj}$ are mapped to $\frac{\rho_{jj}+\mathrm{tr}(\rho)}{d+1}$.
In contrast, for the channel corresponding to uniform sampling from DDBs, the mapping is:
  • The off-diagonal elements $\rho_{jk}$ (where $j\neq k$) are mapped to $\rho_{jk}/(2d)$.
  • The diagonal elements $\rho_{jj}$ are mapped to $\frac{d\times\rho_{jj}+\mathrm{tr}(\rho)}{2d}$.
The inverse reconstruction channel of uniform sampling DDBs is given by
$\mathcal{M}^{-1}(\rho)$ $=2d\Bigg{[}\rho-\frac{d-1}{d}\sum_{k=0}^{d-1}\mbox{tr}(\rho P_{k})P_{k}\Bigg{]}-\frac{\mbox{tr}(\rho)}{d}I.$ (32)
It also includes the linear combination of $\rho$ and $I$, with additional corrections to the diagonal terms.
Appendix B Performance guarantee
The predicted observable is $O$. Denote its traceless part as $O_{0}=O-\mbox{tr}(O)I/d$. If we uniformly sample the DDBs for state $\sigma$ as introduced above, the sampling complexity is linearly dependent on the variance in Eq. (5).
We have
$\mathcal{M}^{-1}(O_{0})=2d\left[O_{0}-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(O_{0}P_{k})P_{k}\right].$ (33)
Denote $\mathcal{M}^{-1}(O_{0})=2d\times o$, where $o=O_{0}-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(O_{0}P_{k})P_{k}$.
By the definition of $o$, we can calculate the relationship between its matrix elements and those of the matrix $O_{0}$:
$&\text{tr}(oP_{k})=\frac{1}{d}\text{tr}(O_{0}P_{k})\\ &\text{tr}(o|j\rangle\langle k|)=\text{tr}(O_{0}|j\rangle\langle k|).\\ \end{aligned}\right.$ (34)
This means that the diagonal elements of the operator $o$ correspond to the reciprocal of the diagonal elements of the operator $O_{0}$ with a denominator of $d$. The operators $o$ and $O_{0}$ share the same non-diagonal elements. The variance for unknown state $\sigma$ is then expressed as follows:
$\lVert O_{0}\rVert_{\sigma}^{2}$ $=\sum_{k=0}^{d-1}\frac{2\text{tr}(\sigma P_{k})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P_{k})+\sum_{0\leq j<k\leq d-1}[\frac{\text{tr}(\sigma P^{+}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P^{+}_{jk})+\frac{\text{tr}(\sigma Q^{+}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})Q^{+}_{jk})]$ (35)
$+\frac{\text{tr}(\sigma P^{-}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P^{-}_{jk})+\frac{\text{tr}(\sigma Q^{-}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})Q^{-}_{jk})]$
$=4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\sigma P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\sigma Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right].$
B.1 Upper bound for the worst case
We know that $\mbox{tr}(\sigma P_{k}),\mbox{tr}(\sigma P_{jk}^{\pm}),\mbox{tr}(\sigma Q_{jk}^{\pm})\leq 1$. To estimate the upper bound of $\lVert O_{0}\rVert_{\sigma}^{2}$, we can roughly conclude
$\lVert O_{0}\rVert_{\sigma}^{2}\leq 4d\sum_{k=0}^{d-1}\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}^{2}(oQ_{jk}^{\pm})\right]).$ (36)
Denote the matrix elements of $o$ as $o_{jk}$, where $j,k\in\{0,\cdots,d-1\}$. We rewrite Eq. (36) as $\lVert O_{0}\rVert_{\sigma}^{2}\leq 2d\times T$, where
$T\doteq$ $\sum_{k=0}^{d-1}2\text{tr}^{2}(oP_{k})+\sum_{0\leq j<k\leq d-1}\left[\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}^{2}(oQ_{jk}^{\pm})\right]$
$=$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\frac{1}{4}\sum_{0\leq j<k\leq d-1}\big{[}(o_{jj}+o_{kk}+o_{jk}+o_{kj})^{2}+(o_{jj}+o_{kk}-o_{jk}-o_{kj})^{2}$
$+(o_{jj}+o_{kk}+io_{jk}-io_{kj})^{2}+(o_{jj}+o_{kk}-io_{jk}+io_{kj})^{2}\big{]}$
$=$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}\big{[}(o_{jj}+o_{kk})^{2}+2o_{jk}o_{kj}\big{]}$
$\leq$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}\big{[}2(o_{jj}^{2}+o_{kk}^{2})+2o_{jk}o_{kj}\big{]}$
$=$ $2d\sum_{k=0}^{d-1}o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}$
$=$ $\frac{2d}{d^{2}}\sum_{k=0}^{d-1}[d\times o_{kk}]^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}.$
When $d\geq 2$, we have $\frac{2d}{d^{2}}\leq 1$. So, we can deduce the upper bound of $T$.
$T\leq\sum_{k=0}^{d-1}[d\times o_{kk}]^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}=\text{tr}(O_{0}^{2}).$
Here we use the relation in Eq. (34).
Thus the upper bound of $\lVert O_{0}\rVert_{\sigma}^{2}$ can be deduced. For each unknown state $\sigma$, the expectation values of $\text{tr}(\sigma P_{k})$, $\text{tr}(\sigma P_{jk}^{\pm})$, and $\text{tr}(\sigma Q_{jk}^{\pm})$ are no bigger than 1. Then we have
$\lVert O_{0}\rVert_{\text{ shadow}}^{2}=\max_{\sigma:\text{ state }}\lVert O_{0}\rVert_{\sigma}^{2}\leq 2dT\leq 2d\times\text{tr}(O_{0}^{2}).$ (37)
Consider the variance in Eq. (35), the first part $4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})$ tends to zero as $d$ increases. As we have $\text{tr}^{2}(oP_{k})=\frac{\mathop{\mathrm{tr}}^{2}(O_{0}|k\rangle\langle k|)}{d^{2}}\leq\mathop{\mathrm{tr}}(O^{2}_{0})/d^{2}$ for all $k$. Thus
$4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})\leq\frac{4\mathop{\mathrm{tr}}(O^{2}_{0})}{d}.$ (38)
Then for bounded-norm observable, the efficiency of sampling complexity is just related to the following part:
$V_{\mbox{diag}}=2d\times\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\sigma P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\sigma Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right].$ (39)
One worst case could happen when the unknown state and the observable are the same as one of the nontrivial DDB states in Eq.(9). For example, $\sigma=P_{01}^{+}$ and $O=P_{01}^{+}$. Then $O_{0}=P_{01}^{+}-I/d$, $o=\frac{d-2}{2d^{2}}(|0\rangle\langle 0|+|1\rangle\langle 1|)+\frac{|0\rangle\langle 1|+|1\rangle\langle 0|}{2}$. Thus $\mathop{\mathrm{tr}}(oP_{jk})>1/2$ and $V_{\mbox{diag}}>d$. Then in this case, the sampling complexity is linear dependent with $d$.
B.2 Average performance analysis
If we sample $|\phi\rangle=U|0\rangle$ with Haar measure, the average state will be
$\int_{U(2^{n})}U|0\rangle\langle 0|U^{{\dagger}}\mathrm{d}\mu(U)=I/2^{n}\;.$ (40)
Thus the variance of state $\sigma=I/d$ exhibits the average performance when the output $\rho$ is randomly and uniformly generated.
When the unknown state is $\sigma=I/d$, we have
$\mathop{\mathrm{tr}}(\sigma P_{k})=\mathop{\mathrm{tr}}(\sigma P^{\pm}_{jk})=\mathop{\mathrm{tr}}(\sigma Q^{\pm}_{jk})=1/d.$ (41)
Take Eq. (41) into Eq. (35), we have
$\lVert O_{0}\rVert_{I/d}^{2}\leq\frac{1}{d}2dT\leq 2\text{tr}(O_{0}^{2}).$ (42)
Thus the average performance is efficient for bounded norm observables. This is constant times of the average performance when we uniformly sample MUB measurements, which is given by $(1+1/2^{n})\text{tr}(O_{0}^{2})$, where $d=2^{n}$. However, the informationally complete $d+1$ MUBs have only been found in prim power dimensions.
B.3 Approximate average state case
If the state $\rho$ is approximately DDB-average, then its deviation from the completely mixed state $I/d$ is small. Specifically, we have
$\left|\text{tr}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d}$
for all snapshots $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$. This implies that $\text{tr}(\rho|\phi\rangle\langle\phi|)\leq\frac{O(\text{poly}(\log d))+1}{d}$. Substituting this into Eq. (35), we can deduce
$\lVert O_{0}\rVert_{\rho}^{2}$ $=4d\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})\cdot\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\rho P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\rho Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right]$ (43)
$\leq O(\text{poly}(\log d)+1)\times 2T$
$\leq O(\text{poly}(\log d))\times\text{tr}(O^{2}_{0}).$
Denote $\epsilon_{1}=\frac{O(\text{poly}(\log d))}{d}$. It is an interesting question to characterize the proportion of randomly chosen states for various levels of deviation $\epsilon_{1}$, such as $\frac{\log d}{d}$, $\frac{\log^{2}d}{d}$, $\frac{\log^{3}d}{d}$, and so forth.
B.4 More exact variance formula
We now present a more concise expression for $V_{\text{diag}}$ in Eq. (39) by substituting the form given in Eq. (27). The variance for $\rho$ and a BN-observable $O$, given by $\lVert O_{0}\rVert_{\sigma}^{2}$, can be expressed as $V_{\text{diag}}$ plus a term that vanishes as $d\to\infty$. Hence, the growth of $\lVert O_{0}\rVert_{\sigma}^{2}$ is determined by the behavior of $V_{\text{diag}}$. If $V_{\text{diag}}$ scales polynomially, then $\lVert O_{0}\rVert_{\sigma}^{2}$ also exhibits polynomial scaling; conversely, if $V_{\text{diag}}$ scales exponentially, $\lVert O_{0}\rVert_{\sigma}^{2}$ will follow an exponential growth pattern as well.
$V_{\text{diag}}$ $=2d\sum_{0\leq j<k\leq d-1}\left[\operatorname{tr}(\sigma P_{jk}^{\pm})\cdot\operatorname{tr}^{2}(oP_{jk}^{\pm})+\operatorname{tr}(\sigma Q_{jk}^{\pm})\cdot\operatorname{tr}^{2}(oQ_{jk}^{\pm})\right]$ (44)
$=\frac{d}{4}\sum_{0\leq j<k\leq d-1}\Big{[}(\sigma_{jj}+\sigma_{kk}+\sigma_{jk}+\sigma_{kj})(o_{jj}+o_{kk}+o_{jk}+o_{kj})^{2}$
$\quad+(\sigma_{jj}+\sigma_{kk}-\sigma_{jk}-\sigma_{kj})(o_{jj}+o_{kk}-o_{jk}-o_{kj})^{2}$
$\quad+(\sigma_{jj}+\sigma_{kk}+i\sigma_{jk}-i\sigma_{kj})(o_{jj}+o_{kk}+io_{jk}-io_{kj})^{2}$
$\quad+(\sigma_{jj}+\sigma_{kk}-i\sigma_{jk}+i\sigma_{kj})(o_{jj}+o_{kk}-io_{jk}+io_{kj})^{2}\Big{]}$
$=\frac{d}{4}\sum_{0\leq j<k\leq d-1}\Big{[}4(\sigma_{jj}+\sigma_{kk})(o_{jj}+o_{kk})^{2}$
$\quad+4(\sigma_{jk}+\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}+o_{kj})$
$\quad-4(\sigma_{jk}-\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}-o_{kj})$
$\quad+8(\sigma_{jj}+\sigma_{kk})o_{jk}o_{kj}\Big{]}$
$=d\sum_{0\leq j<k\leq d-1}[(\sigma_{jj}+\sigma_{kk})(o_{jj}+o_{kk})^{2}+(\sigma_{jk}+\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}+o_{kj})-(\sigma_{jk}-\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}-o_{kj})$
$\quad+2(\sigma_{jj}+\sigma_{kk})o_{jk}o_{kj}].$
Now, consider the case where $O$ is an off-diagonal observable, meaning $O_{jj}=o_{jj}=0$ for all $j=0,\ldots,d-1$. Under this condition, the expression simplifies to:
$V_{\text{diag}}=2d\sum_{0\leq j<k\leq d-1}(\sigma_{jj}+\sigma_{kk})o_{jk}o_{kj}.$
B.5 Proof of lemma 2: Efficient Clifford circuit construction for stabilizer states
Analysis: By Eq. (42) of [23], any stabilizer state can be expressed as
$|\Psi\rangle=\omega U_{C}U_{H}|s\rangle,$
where $U_{C}$ and $U_{H}$ are $C$-type and $H$-type Clifford operators, $s\in\{0,1\}^{n}$ is a basis vector, and $\omega$ is a complex phase factor. The $C$-type Clifford operators are those that can be decomposed into gates from the set $\{S,\text{CZ},\text{CX}\}$, while the $H$-type Clifford operators only consist of Hadamard gates. Furthermore, a polynomial-time algorithm is provided to efficiently obtain this CH-form of the stabilizer state using the stabilizer tableaux representation.
Consequently, the target Clifford circuit $T$ can be decomposed into $U_{C}^{\dagger}$, followed by a sequence of SWAP operations (each equivalent to three CX gates). ∎
B.6 Proof of theorem 3: Efficient estimation of expectation values using DDB-ST
Proof.
We categorize the cases based on the number of non-zero coefficients present in $|\Psi\rangle\langle\Psi|$.
For $n-\log(\operatorname{poly}(n))\leq r\leq n$, the stabilizer state $|\Psi\rangle$ is approximately DDB-average. Thus, $O(\operatorname{poly}(n))$ samples and computational resources are sufficient for efficient estimation using $n$-qubit DDB-ST (Lemma 1).
Next, for $0\leq r\leq\log(\operatorname{poly}(n))$, the matrix $|\Psi\rangle\langle\Psi|$ is sparse, containing at most $4^{r}$ non-zero coefficients, making direct computation of $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ efficient.
For a general $r$ and given $|\Psi\rangle$, we can efficiently find a Clifford circuit $U$ such that $|\Psi\rangle=U|j\rangle$ using the Gottesman-Knill Theorem. By Lemma 2, we can further identify a unitary operation $T$ such that $|\Phi\rangle=T|\Psi\rangle=TU|j\rangle$, where
$|\Phi\rangle$ $=|0\rangle^{\otimes(n-r)}\otimes\frac{1}{\sqrt{2^{r}}}\sum_{k\in\{0,1\}^{r}}\omega^{q^{\prime}(k)}|k\rangle$ (45)
$=|0\rangle^{\otimes(n-r)}\otimes|\Phi_{r}\rangle.$
In other words, we obtain a new stabilizer state $|\Phi\rangle$ where only the first $2^{r}$ components are non-zero.
The target expectation value can be rewritten as:
$\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ $=\mathrm{tr}(T|\Psi\rangle\langle\Psi|T^{\dagger}TOT^{\dagger})$ (46)
$=\mathrm{tr}(|\Phi\rangle\langle\Phi|\times TOT^{\dagger}).$
Given that the matrix form of $|\Phi\rangle\langle\Phi|$ is
$\begin{bmatrix}|\Phi_{r}\rangle\langle\Phi_{r}|&\mathbf{0}\\ \mathbf{0}&\mathbf{0}_{2^{n-r}\times 2^{n-r}}\end{bmatrix},$
we only need to consider the leading $2^{r}\times 2^{r}$ submatrix of $TOT^{\dagger}$, denoted as $[TOT^{\dagger}]_{2^{r}\times 2^{r}}$.
Thus, the computation of $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ simplifies to:
$\mathrm{tr}(|\Phi_{r}\rangle\langle\Phi_{r}|\times[TOT^{\dagger}]_{2^{r}\times 2^{r}}).$
Since $\mathop{\mathrm{tr}}[(TOT^{\dagger})^{2}]=\mathop{\mathrm{tr}}(O^{2})$ and $[TOT^{\dagger}]_{2^{r}\times 2^{r}}$ is a submatrix, we can conclude that $\mathop{\mathrm{tr}}\left([TOT^{\dagger}]_{2^{r}\times 2^{r}}^{2}\right)$ is also bounded by the same constant as $\mathop{\mathrm{tr}}(O^{2})$.
Furthermore, the $r$-qubit stabilizer state $|\Phi_{r}\rangle$ is an $r$-qubit DDB-average state by Property 1. Therefore, we can apply $r$-qubit DDB-ST to estimate $\mathrm{tr}(|\Phi_{r}\rangle\langle\Phi_{r}|\times[TOT^{\dagger}]_{2^{r}\times 2^{r}})$. In the post-processing step, we will compute polynomial expectation values of the form:
$\mathrm{tr}(\mathcal{M}^{-1}(|\phi_{r}\rangle\langle\phi_{r}|)\times[TOT^{\dagger}]_{2^{r}\times 2^{r}}),$
where $|\phi_{r}\rangle$ is an $r$-qubit DDB state, and $\mathcal{M}^{-1}$ is the inverse channel in Eq. (11) with dimension $d=2^{k}$. Let $\tau_{r}=\mathcal{M}^{-1}(|\phi_{r}\rangle\langle\phi_{r}|)$. It is evident that $\tau_{r}$ contains at most four non-zero matrix elements.
Even though $[TOT^{\dagger}]_{2^{r}\times 2^{r}}$ is a matrix of size $2^{r}\times 2^{r}$, its explicit form is not required. Let $\tau=(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}$, then we have:
$\mathrm{tr}(\mathcal{M}^{-1}(|\phi_{r}\rangle\langle\phi_{r}|)\times[TOT^{\dagger}]_{2^{r}\times 2^{r}})$ $=\mathrm{tr}(\tau\times TOT^{\dagger})$ (47)
$=\mathrm{tr}(T^{\dagger}\tau T\times O).$
Since the $n$-qubit operator $\tau$ has at most four non-zero elements, $T^{\dagger}\tau T$ can be efficiently computed using the Gottesman-Knill theorem with complexity $\text{poly}(n)$. Moreover, $T$ can be decomposed into gates from the set $\{S,\text{CZ},\text{CX}\}$, ensuring that $T^{\dagger}\tau T$ also contains at most four non-zero elements. As when $C$-type Clifford operators act on a computational basis state, the resulting state remains in the computational basis, multiplied by a global phase.
Thus, in each post-processing step of $r$-qubit DDB-ST, the computational complexity is $\text{poly}(n)\times\text{constant}$. The first $\text{poly}(n)$ term accounts for the complexity of obtaining $T^{\dagger}\tau T$, while the constant factor corresponds to the calculation of $\mathrm{tr}(T^{\dagger}\tau T\times O)$. With polynomial samples of DDB-ST, the total computational complexity for estimating $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ remains polynomial.
Therefore, the estimation of $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ in all cases requires only polynomial samples and computations.