### v5

# Quantum Advantage via Efficient Post-processing on Qudit Classical Shadow Tomography

Abstract
Computing inner products of the form $\operatorname{tr}(AB)$, where $A$ is a $d$-dimensional density matrix (with $\operatorname{tr}(A)=1$, $A\geq 0$) and $B$ is a bounded-norm observable (Hermitian with $\operatorname{tr}(B^{2})\leq O(\mathrm{poly}(\log d))$ and $\operatorname{tr}(B)$ known), is fundamental across quantum science and artificial intelligence. Classically, both computing and storing such inner products require $O(d^{2})$ resources, which rapidly becomes prohibitive as $d$ grows exponentially. In this work, we introduce a quantum approach based on qudit classical shadow tomography, significantly reducing computational complexity from $O(d^{2})$ down to $O(\mathrm{poly}(\log d))$ in typical cases and at least to $O(d~{}\text{poly}(\log d))$ in the worst case. Specifically, for $n$-qubit systems (with $n$ being the number of qubit and $d=2^{n}$), our method guarantees efficient estimation of $\operatorname{tr}(\rho O)$ for any known stabilizer state $\rho$ and arbitrary bounded-norm observable $O$, using polynomial computational resources. Crucially, it ensures constant-time classical post-processing per measurement and supports qubit and qudit platforms. Moreover, classical storage complexity of $A$ reduces from $O(d^{2})$ to $O(m\log d)$, where the sample complexity $m$ is typically exponentially smaller than $d^{2}$. Our results establish a practical and modular quantum subroutine, enabling scalable quantum advantages in tasks involving high-dimensional data analysis and processing.
Introduction—Computing $\text{tr}(AB)$ for two known $d$-dimensional Hermitian operators is fundamental in quantum science and artificial intelligence. When $A$ and $B$ lack structure (e.g., sparsity or low rank), the classical computational and storage costs scale as $O(d^{2})$, which becomes prohibitive for exponentially large $d$. In quantum settings, similar quantities arise: the expectation value $\text{tr}(\rho O)$ predicts the outcome of measuring an observable $O$ on an unknown quantum state $\rho$. Such predictions are central to quantum information processing, quantum simulation, and quantum chemistry. However, when $\rho$ is unknown and the dimension is exponential, estimating $\text{tr}(\rho O)$ efficiently becomes significantly more challenging.
Fortunately, classical shadow tomography based on random Clifford measurements (Clifford-ST) provides a scalable approach for efficiently estimating $\text{tr}(\rho O)$, when $\rho$ is an unknown $n$-qubit state and $O$ is a bounded-norm (BN) observable satisfying $\text{tr}(O^{2})\leq O(\text{poly}(n))$ Huang et al. (2020), and given that $\text{tr}(O)$ is known. It yields a substantial quantum advantage in experimental learning tasks, exponentially reducing the sample complexity relative to classical methods Huang et al. (2022).
Quantum advantage plays a central role in quantum computing, with representative breakthroughs including the Deutsch–Jozsa algorithm Deutsch and Jozsa (1992), Shor’s factoring algorithm Shor (1994), Grover’s search algorith Grover (1996), and the HHL algorithm Harrow et al. (2009). As large-scale matrix operations become increasingly common in quantum science and artificial intelligence, it is natural to ask whether shadow-based methods—originally designed for quantum state tomography—can be repurposed as scalable quantum subroutines algorithm for computing quantities such as $\text{tr}(AB)$, delivering broader advantages in both sampling and computational complexity.
However, several challenges limit existing shadow estimation protocols. In Clifford-ST, each single-shot measurement yields an estimator $\tilde{\rho}$, and post-processing involves computing $\text{tr}(\tilde{\rho}O)$, which can be computationally expensive when $\tilde{\rho}$ is dense and $O$ lacks efficient stabilizer expression. In the worst case, the post-processing cost can become exponential. Recent work has extended Clifford-ST from qubit systems ($d=2^{n}$) to qudits of odd prime power dimensions Mao et al. (2024). Meanwhile, MUB-based protocols have been proposed for BN observables Wang and Cui (2024), but the existence of a complete set of $d+1$ MUB remains unresolved for general $d$ Horodecki et al. (2022), limiting their applicability to general dimension $d$. On near-term quantum devices, implementing random Clifford circuits with $O(n^{2})$ decomposed gates remains experimentally challenging. A workaround on optical quantum platforms is to project onto random stabilizer states, which is mathematically equivalent but requires classical preprocessing with complexity $O(n^{3}2^{n})$ struchalin2021experimental .
These challenges highlight the need for alternative shadow estimation protocols that simultaneously reduce classical post-processing overhead, support general dimensionality, and remain experimentally feasible.
In this work, we introduce DDB-ST (Dense Dual Bases Classical Shadow Tomography), a new protocol based on randomized projective measurements over $2d$ dense dual bases. Our scheme supports arbitrary finite dimension $d$ and ensures constant-time classical post-processing per measurement—independent of the quantum state $\rho$ or observable $O$. Consequently, the total classical computational cost scales linearly with sample complexity. For BN observables, the worst-case sample complexity is $O(d~{}\text{poly}(\log d))$, while in the average case it achieves exponential speedup over classical methods. In particular, for any $n$-qubit stabilizer state $\rho$ and BN observable $O$, our method estimates $\text{tr}(\rho O)$ using only $\text{poly}(n)$ resources.
Beyond its role in quantum state estimation, we further explore DDB-ST as a modular subroutine in machine learning and quantum simulation pipelines, where it significantly reduces memory overhead. In particular, by replacing full matrix representations with compact classical shadows, DDB-ST compresses the storage cost of a density matrix $A$ from $O(d^{2})$ to $O(m\log d)$, where $m$ is the number of measurement samples. Moreover, DDB-ST can be directly applied to efficiently estimate $\text{tr}(AB)$, where $A$ is a quantum state and $B$ a known observable. This enables simultaneous reductions in both classical storage and computational complexity, which are central bottlenecks in quantum science, optimization, and machine learning. These features position DDB-ST as a scalable, hardware-efficient, and broadly applicable primitive for quantum-enhanced data processing.
Preliminaries and framework—Classical shadow tomography, as introduced by Huang, Kueng, and Preskill Huang et al. (2020), provides a method for efficiently predicting properties of an unknown quantum state using randomized measurements.
The process involves randomly sampling a unitary transformation $U_{k}$ with probability $p_{k}$ from an informationally complete ensemble $\mathcal{U}=\{U_{k}\}$, evolving the state $\rho\to U_{k}\rho U_{k}^{\dagger}$, and performing projective measurements in the computational basis. It corresponds to a quantum process
$\mathcal{M}(\rho)=\sum_{k}p_{k}\mathop{\mathrm{tr}}(U_{k}\rho U_{k}^{{\dagger}}|j\rangle\langle j|)U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}.$ (1)
Each experiment yields a single estimator of the unknown state $\rho$:
$\tilde{\rho}=\mathcal{M}^{-1}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k}).$ (2)
The estimation under observable $O$ by one measurement is given by
$\mathop{\mathrm{tr}}(\mathcal{M}^{-1}(U_{k}^{{\dagger}}|j\rangle\langle j|U_{k})~{}O).$ (3)
Let $\langle o_{k}\rangle$ as the final estimation of $\mathop{\mathrm{tr}}(\rho O_{k})$. Since the quantum state collapses upon measurement, multiple copies of the unknown state should be prepared. Given exponentially large number of observables $\{O_{k}\}_{k=1}^{L}$, to achieve accuracy $\epsilon$ and confidence level $1-\sigma$, we can ensure
$P_{r}(|\langle o_{k}\rangle-\mathop{\mathrm{tr}}(\rho O_{k})|<\epsilon)\geq 1-\sigma$ (4)
with the following sample complexity:
$N=O\left(\frac{\log\frac{L}{\sigma}}{\epsilon^{2}}\right)\max_{1\leq i\leq L}\left\|O_{i}-\frac{\text{tr}(O_{i})}{d}I\right\|_{\mbox{shadow}}^{2},$ (5)
where $\|\cdot\|_{\mbox{shadow}}^{2}=\max_{\sigma:\text{ state }}\|\cdot\|_{\sigma}^{2}$ denotes the shadow norm associated with $\mathcal{U}$ and observables $\{O_{k}\}$. It represents the worst-case sample complexity, while $\|\cdot\|_{\sigma}^{2}$ is related to the sample complexity for state $\sigma$. Denote $O_{0}=O-\mathop{\mathrm{tr}}(O)I/d$, $\lVert O_{0}\rVert_{\sigma}^{2}$ is defined as
$\mathbb{E}_{U\sim\mathcal{U}}\sum_{b\in\{0,1\}^{n}}\langle b|U\sigma U^{\dagger}|b\rangle\cdot\langle b|U\mathcal{M}^{-1}(O_{0})U^{\dagger}|b\rangle^{2}.$ (6)
If the ensemble $\mathcal{U}$ consists of all $n$-qubit Clifford circuits, the inverse channel admits a closed-form:
$\mathcal{M}^{-1}(X)=(2^{n}+1)X-\operatorname{tr}(X)\,I.$
The variance of Clifford-ST satisfies $\|\cdot\|_{\text{shadow}}\leq 3\operatorname{tr}(O^{2})$, ensuring that BN observables—those with $\operatorname{tr}(O^{2})\leq\mathrm{poly}(n)$—can be estimated using $O(\mathrm{poly}(n))$ samples.
Clifford-ST can be viewed as a quantum algorithm for approximating $\mathop{\mathrm{tr}}(\rho O)$ to a target precision. For each measurement, recording $U_{k}$ and outcome $j$, the post-processing computes
$\text{tr}(\tilde{\rho}O)=(2^{n}+1)\,\mathop{\mathrm{tr}}(U_{k}^{\dagger}|j\rangle\langle j|U_{k}O)-\,\mathop{\mathrm{tr}}(O),$ (7)
and the final estimate is given by $\mathop{\mathrm{tr}}(\rho O)\approx\sum_{k=1}^{T(n)}\frac{\text{tr}(\tilde{\rho}O)}{T(n)}$, where $T(n)$ is the sample complexity (see Eq. (5)), which is $O(\mathrm{poly}(n))$ for BN observables. While statistical techniques such as median-of-means improve robustness, they do not reduce the scaling of overall complexity.
The computational cost for Eq. (7) depends on the structure of $\tilde{\rho}$ and $O$. In Clifford-ST, the projected states, or snapshots, lie in the set
$\mathcal{S}_{\text{Clifford}}=\left\{U_{k}^{\dagger}|j\rangle:U_{k}\in\mathcal{U}_{\text{Clifford}},\ j=0,\dots,d-1\right\},$
which contains $O(2^{n^{2}})$ $n$-qubit stabilizer states. Due to the classical complexity of computing $\mathop{\mathrm{tr}}(AB)$, evaluating $\mathop{\mathrm{tr}}(\tilde{\rho}O)$ often requires $O(2^{2n})$ operations when both matrices are general.
There are two favorable cases where the post-processing of Clifford-ST becomes efficient:
  • Sparsity of measurement states: If all collapsed $\{U_{k}^{\dagger}|j\rangle\}$ are sparse with $O(\mathrm{poly}(n))$ nonzero amplitudes, then for arbitrary $O$, the trace can be computed in polynomial time. However, such sparse states constitute only a small fraction of $\mathcal{S}_{\text{Clifford}}$.
  • Stabilizer-decomposable observables: If the observable admits a decomposition into a polynomial number of stabilizer projectors,
    $O=\sum_{l=1}^{\mathrm{poly}(n)}c_{l}\,|\phi_{l}\rangle\langle\psi_{l}|,\quad\text{with }|\phi_{l}\rangle,|\psi_{l}\rangle\in\mathcal{S}_{\text{Clifford}},$
    then $\mathop{\mathrm{tr}}(\tilde{\rho}O)$ can be efficiently computed using the Gottesman–Knill theorem.
Unfortunately, such observables are low-rank, form a measure-zero subset of all BN observables. As a result, despite the favorable sampling complexity of Clifford-ST, its classical post-processing cost can be exponential in the worst case.
To avoid this overhead, we modify the set of projected states $\mathcal{S}=\{U_{k}^{\dagger}|j\rangle\langle j|U_{k}\}$ such that, for any $d$-dimensional observable $O$, the computational cost of evaluating $\mathop{\mathrm{tr}}(\tilde{\rho}O)$ remains constant.
Snapshots with dense dual basis states— We define the following states:
$\begin{cases}|\phi_{jk}^{\pm}\rangle=\frac{1}{\sqrt{2}}(|j\rangle\pm|k\rangle),\\ |\psi_{jk}^{\pm}\rangle=\frac{1}{\sqrt{2}}(|j\rangle\pm i|k\rangle).\end{cases}$ (8)
The new collection of snapshots comprises a total of $2d^{2}-d$ elements:
$\mathcal{S}_{\text{DDB}}=\{P_{t}=|t\rangle\langle t|,~{}~{}t=0,\ldots,d-1;\\ P_{jk}^{\pm}=|\phi_{jk}^{\pm}\rangle\langle\phi_{jk}^{\pm}|,Q_{jk}^{\pm}=|\psi_{jk}^{\pm}\rangle\langle\psi_{jk}^{\pm}|;~{}~{}0\leq j<k\leq d-1\}.$ (9)
These rank-1 projectors are informationally complete, as their linear span covers the entire space of $d\times d$ Hermitian operators $\mathbb{M}_{d}(\mathbb{C})$. This property makes such sets suitable for classical shadow tomography.
An efficient algorithm with $O(\log d)$ iterations has been developed to construct a unitary ensemble $\mathcal{U}_{\text{DDB}}=\{U_{j}\}_{j=1}^{f(d)}$, containing the minimal number of elements required to span all rank-1 projectors in Eq. (9Wang et al. (2024). Here, $f(d)=2d$ when $d$ is odd, and $f(d)=2d-1$ when $d$ is even. Each orthonormal basis $\{U_{j}|k\rangle:k=0,\dots,d-1\}$ is referred to as a Dense Dual Basis (DDB).
In $n$-qubit systems, each DDB circuit consists of a single Hadamard gate (optionally followed by a phase gate $S$) and a permutation gate, and can be decomposed into $O(n^{4})$ elementary gates. Although this gate count is higher than that of Clifford circuits, random projections onto DDB states yield exponentially improved classical pre-processing efficiency compared to projections onto stabilizer states struchalin2021experimental , as detailed in the Supplementary Material.
Moreover, DDB-ST generalizes naturally to arbitrary finite dimensions. On $n$-qubit systems, the total number of DDB states is $O(2^{2n})$, significantly fewer than the $O(2^{n^{2}})$ stabilizer states, allowing for a biased but efficient sampling strategy. In the following, we investigate the explicit reconstruction channel of DDB-ST, along with its sample complexity and classical computational cost.
Theorem 1 (Reconstruction channel for DDB-ST).
Let $\rho=\sum_{j,k=0}^{d-1}\rho_{jk}|j\rangle\langle k|$, and define $P_{k}=|k\rangle\langle k|$ as the projectors onto the computational basis.
  • For odd dimensions $d$, each unitary $U_{j}$ in the ensemble $\mathcal{U}_{\text{DDB}}$ is sampled uniformly with probability $1/(2d)$.
  • For even dimensions $d$, the ensemble $\mathcal{U}_{\text{DDB}}$ is sampled, where the identity $I$ (corresponding to the computational basis) is selected with probability $2/(2d)$, and each remaining $U_{j}$ with probability $1/(2d)$.
The corresponding quantum channel $\mathcal{M}$ takes the form:
$\mathcal{M}(\rho)=\frac{1}{2d}\left[\rho+\operatorname{tr}(\rho)\,I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}P_{k}\right].$ (10)
Its inverse reconstruction channel $\mathcal{M}^{-1}$ is given by:
$\mathcal{M}^{-1}(\rho)=2d\left[\rho-\frac{d-1}{d}\sum_{k=0}^{d-1}\operatorname{tr}(\rho P_{k})\,P_{k}\right]-\frac{\operatorname{tr}(\rho)}{d}\,I.$ (11)
All detailed derivations and proofs are provided in the Supplementary Material.
Property 1.
For any estimated state and any $d$-dimensional Hermitian observable $O$ with known trace, the classical post-processing complexity of Eq. (3) in a single-shot measurement of DDB-ST is five, i.e., constant in $d$, regardless of the structure of $O$.
Theorem 2 (Performance Guarantee).
In a $d$-dimensional Hilbert space, when using DDB-ST to predict the expectation value of any observable $O$, the worst-case variance for each quantum state $\sigma$ is bounded by:
$\lVert O_{0}\rVert_{\text{shadow}}^{2}=\max_{\sigma:\text{state}}\lVert O_{0}\rVert_{\sigma}^{2}\leq 2d\,\operatorname{tr}(O_{0}^{2}),$ (12)
where $O_{0}=O-\frac{\operatorname{tr}(O)}{d}I$. If the unknown state is sampled randomly according to the Haar measure, the average variance is bounded by:
$\lVert O_{0}\rVert_{I/d}^{2}\leq 2\,\operatorname{tr}(O_{0}^{2}).$ (13)
Definition 1 (Approximately DDB-Average State).
A state $\rho$ is called approximately DDB-average if it satisfies the following condition:
$\left|\text{tr}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d},$ (14)
for all $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$.
The maximally mixed state $\rho=I/d$ satisfies $\mathop{\mathrm{tr}}(\rho|\phi\rangle\langle\phi|)=1/d$ for all $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$, and is thus exactly DDB-average.
We performed numerical simulations to estimate the proportion of approximately DDB-average states (Fig. 1). For each dimension $d=2^{2},\ldots,2^{8}$, we generated 1000 Haar-random pure states $\rho$, labeling a state as approximately DDB-average if $\max_{|\phi\rangle\in\mathcal{S}_{\text{DDB}}}\left|\mathop{\mathrm{tr}}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{s}{2^{n}}.$ Fig. 1 reports the fraction of such states for thresholds $s=4$ (green), $s=2n$ (orange), and $s=n^{2}$ (blue). The proportion rapidly increases with $s$, reaching 100% when $s=n^{2}$.
Interestingly, MUB-based classical shadow tomography has average variance $(1+1/2^{n})\mathop{\mathrm{tr}}(O_{0}^{2})$, roughly half that of DDB-ST. For small $n$, thresholds $s=n$ and $s=2n$ suffice to capture most Haar-random states as approximately MUB- or DDB-average, respectively.
Moreover, we generated random mixed states via the Hilbert–Schmidt measure using Qiskit. For $n=2$ to 7 and threshold $s=2$, all sampled states satisfied the DDB-average condition, suggesting that mixed states more readily meet this criterion.
Figure 1: Proportion of approximately DDB-average states.
Lemma 1.
For any BN-observable $O$ and an approximately DDB-average state $\rho$, the variance of DDB-ST satisfies: $\lVert O_{0}\rVert_{\rho}^{2}\leq O(\operatorname{poly}(\log d))\cdot\operatorname{tr}(O_{0}^{2}).$ As a result, the sample and computational complexity for estimating $\operatorname{tr}(\rho O)$ using DDB-ST is $O(\operatorname{poly}(\log d))$.
Complexity comparison—We systematically compare sampling complexity, and classical post-processing complexity across classical methods, Clifford-ST, and DDB-ST.
Classically, computing $\operatorname{tr}(AB)$ for a general $d$-dimensional density matrix $A$ and BN observable $B$ requires $O(d^{2})$ time and storage.
For shadow-based approaches, Clifford-ST achieves a sample complexity of $O(\mathrm{poly}(n))$ when estimating $\operatorname{tr}(\rho O)$ for any $n$-qubit state $\rho$ and BN observable $O$. However, the classical post-processing cost is between $O(\mathrm{poly}(n)\cdot p(n))$ and $O(\mathrm{poly}(n)\cdot P(n))$, where $p(n)$ and $P(n)$ denote the best and worst cases in all measurements, respectively. In unfavorable scenario, the cost $p(n)$ can scale as $O(4^{n})$.
In contrast, DDB-ST achieves sample complexity $O(d\cdot\mathrm{poly}(\log d))$ in the worst case and $O(\mathrm{poly}(\log d))$ on average (see Eqs. (12) and (13)), with constant-time classical post-processing per measurement, independent of $\rho$ and $O$. The total classical computational complexity of estimating $\operatorname{tr}(\rho O)$ is
$\mathcal{N}_{\text{DDB}}\leq O\left(\frac{\log\frac{1}{\sigma}}{\epsilon^{2}}\right)\cdot\lVert O_{0}\rVert_{\rho}^{2}\cdot\text{constant},$ (15)
where $\epsilon$ is the desired precision and $\sigma$ is the failure probability (which controls the confidence level).
This implies exponential speedup whenever $\lVert O_{0}\rVert_{\rho}^{2}=O(\mathrm{poly}(\log d))$, and guarantees at least quadratic improvement for all $\rho$ in the worst case due to the bound $\lVert O_{0}\rVert_{\text{shadow}}^{2}\leq 2d\operatorname{tr}(O_{0}^{2})$. Since $\operatorname{tr}(O_{0}^{2})\leq\operatorname{tr}(O^{2})$, when $\operatorname{tr}(O^{2})$ is constant, the worst-case sample complexity scales as $O(d)$. On average, polynomial complexity can be achieved when $\rho$ is randomly chosen according to Haar measure.
Any $n$-qubit stabilizer state Dehaene and De Moor (2003); Nest (2008) is expressed by
$|\Psi\rangle=\frac{1}{\sqrt{2^{r}}}\sum_{k\in A}\omega^{q(k)}|k\rangle,$ (16)
where $\omega^{q(k)}$ is in $\{1,-1,i,-i\}$, $r=0,1,\cdots,n$, and $A=\{Ru+t\mid u\in\mathbb{Z}_{2}^{r}\},$ with $R$ being a nonsingular $n\times r$ matrix, and $t\in\mathbb{Z}_{2}^{N}$ a fixed vector.
Property 2.
For any stabilizer state $|\Psi\rangle$ expressed in Eq. (32), we have
$\max_{|\phi\rangle\in\mathcal{S}_{\text{DDB}}}\text{tr}(|\Psi\rangle\langle\Psi|\phi\rangle\langle\phi|)\leq\frac{1}{2^{r}}.$ (17)
Theorem 3 (Informal).
For any $n$-qubit stabilizer state $|\Psi\rangle$ in Eq. (32) and any BN observable $O$, the expectation value $\operatorname{tr}(|\Psi\rangle\langle\Psi|O)$ can be efficiently estimated using DDB-ST with additive error $\epsilon+\operatorname{tr}(O^{2})/\sqrt{2^{r}}$, requiring $O(\mathrm{poly}(n))$ samples and post-processing time. When $O$ is traceless, the estimation error can be made arbitrarily small. For small $r$, the expectation value can can also be computed accurately and efficiently via direct classical methods.
This result does not contradict the Gottesman-Knill theorem Gottesman (1997), which allows exact computation of $\operatorname{tr}(\rho O)$ when both $\rho$ and $O$ are of stabilizer type. In contrast, our method extends efficient estimation to arbitrary BN observables $O$, even when $O$ lacks an efficient stabilizer decomposition. The cost is an additive error $\epsilon+\operatorname{tr}(O^{2})/\sqrt{2^{r}}$, where the second term decays exponentially with $r$.
Beyond stabilizer states, DDB-ST also provides exponential speedup for a broader class of approximately DDB-average states. Consider the mixed state: $\rho_{A}=\frac{I}{d}+\sum_{j\neq k}\rho_{jk}|j\rangle\langle k|$, with $|\rho_{jk}|<\frac{\mathrm{poly}(\log d)}{d}.$ For any BN-observable $O$, we have $\left|\operatorname{tr}(\rho_{A}O)-\frac{\operatorname{tr}(O)}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d}\sum_{j\neq k}|O_{jk}|\leq\text{poly}(\log d)\,\sqrt{\operatorname{tr}(O^{2})}$. DDB-ST estimates this quantity with $\epsilon$-accuracy using $O(\mathrm{poly}(\log d))$ computational resources. In contrast, $\rho_{A}$ and $O$ each involve $O(d^{2})$ variables, making it exponentially hard to perform the same estimation using purely classical computations.
A relevant physical example is the depolarizing channel, a fundamental noise model in quantum information theory, defined as $\mathcal{D}_{p}(\rho)=(1-p)\rho+p\cdot\frac{I}{d}$ where $p\in[0,1]$. When $p$ is large, the output state $\mathcal{D}_{p}(\rho)$ resembles $\rho_{A}$, making DDB-ST an efficient tool for estimating $\operatorname{tr}[\mathcal{D}_{p}(\rho)O]$ in noisy quantum systems for BN observables.
Applications beyond state learning— Beyond efficiently estimating the properties of quantum states, DDB-ST could serves as a general-purpose quantum subroutine for evaluating trace expressions of the form $\operatorname{tr}(AB)$.
Such evaluations arise naturally in quantum algorithms where the output is a quantum state $|x\rangle$, typically encoded in the amplitudes of a superposition and inaccessible via a single measurement. We assume that generating $|x\rangle$ through quantum computation is not slower than classical counterparts, though many quantum algorithms seek exponential speedup at this stage. For instance, in the HHL algorithm Harrow et al. (2009), or its extensions in quantum machine learning Biamonte et al. (2017), the final state encodes the solution to $A\vec{x}=\vec{b}$ as $|x\rangle$, and estimating $\operatorname{tr}(|x\rangle\langle x|O_{k})$ for observables $\{O_{k}\}$ yields interpretable outputs for special application. Efficiently performing this estimation is essential to realizing end-to-end quantum advantage, as shown in Fig. 2.
Figure 2: Classical vs quantum estimation pipeline: enabling advantage with DDB-ST. DDB-ST offers exponential or quadratic speedup in the estimation stage, bridging the gap between quantum state outputs and downstream applications in AI, optimization, and scientific computing Rao et al. (1973); Horn and Johnson (2012). Although a quantum state $\rho$ has limited lifetime, it could be compactly stored as polynomial-size classical data via well-designed random measurements by classical shadow tomography.
Conclusion and discussion—In this work, we introduce DDB-ST, a shadow tomography framework based on randomized projective measurements over $2d$ DDBs. This method enables efficient estimation of $\operatorname{tr}(AB)$, where $A$ is a $d$-dimensional density matrix and $B$ is a BN observable. Compared to classical methods that require $O(d^{2})$ resources for computation and storage, our method achieves $O(d\cdot\mathrm{poly}(\log d))$ complexity in the worst case and $O(\mathrm{poly}(\log d))$ in typical cases. Unlike Clifford-ST, which may suffer from exponential post-processing cost in general, DDB-ST ensures constant-time post-processing per measurement regardless of $A$ or $B$, and applies to both qubit and qudit systems. Notably, our framework supports efficient estimation for arbitrary BN observables and $n$-qubit stabilizer states, further demonstrating its practical versatility.
There are several promising directions for future research. First, expanding the class of matrices $A$ that allow for polynomial sampling complexity is crucial. This could involve biased sampling, imposing additional constraints on observables, or exploring sparse alternative snapshot sets. In $n$-qubit systems, DDB and MUB measurements represent two extremes of Clifford measurements. Nontrivial DDB and MUB states correspond to stabilizer states in Eq. (32) with $r=1$ and $r=n$ respectively. Both DDB and MUB snapshots contain $O(2^{2n})$ elements and exhibit similar variance properties. Including more snapshots with only polynomial nonzero amplitudes may reduce the worst-case sample complexity.
Second, while our work primarily focuses on the efficient computation of $\mathop{\mathrm{tr}}(AB)$, extending these techniques to nonlinear properties, such as purity and entropy McGinley and Fava (2023), could significantly broaden their range of applications.
Third, many current quantum algorithms focus on achieving exponential speedups in Stage One (Fig. 2) using fixed quantum circuits. Incorporating measurements or randomized quantum circuits could unlock new potential for quantum advantage. In randomized measurement protocols, we design the unitary $\mathcal{U}=\{U_{j}\}$, and nature replies with a measurement outcome $|k\rangle$—a stochastic echo from which we infer the properties of the quantum system. This interplay between controlled randomness and measurement-induced information reflects a deeper structure: even in non-determinism, nature leaves behind a traceable signature signature that is beneficial for learning.
Finally, improving the robustness and accuracy of the scheme in noisy environments is crucial Chen et al. (2021); Koh and Grewal (2022); Wu and Koh (2024). Since our primary goal is to estimate expectation values rather than fully reconstruct quantum states, the task would require less extensive error correction than typical quantum computations.
Acknowledgements— This work received support from the National Natural Science Foundation of China through Grants No. 62001260 and No. 42330707, and from the Beijing Natural Science Foundation under Grant No. Z220002.
Appendix: Supplementary Material
Appendix A Proof of theorem 1: Calculation of the reconstruction channel
We may express $\rho$ with the following form
$\rho=\sum_{j,k=0}^{d-1}\rho_{jk}|j\rangle\langle k|.$ (18)
The DDB unitary ensemble is $\{U_{j}\}_{j=1}^{f(d)}$. For even $d$, $f(d)=2d-1$. The first DDB corresponds to the computational basis, while the remaining bases are constructed in dual pairs, denoted as $\{|\phi_{jk}^{\pm}\rangle:(j,k)\in\mathbb{T}\}$ and $\{|\psi_{jk}^{\pm}\rangle:(j,k)\in\mathbb{T}\}$, where $\mathbb{T}$ is a partition of $\{0,1,\dots,d-1\}$ into distinct pairs with no repeated elements. For example, when $d=4$, one possible partition is $\mathbb{T}_{1}=\{(0,1),(2,3)\}$. A minimum of $d-1$ such partitions is needed to cover all pairs $(j,k)$ where $10\leq j<k\leq d-$ 1.
For odd $d$, $f(d)=2d$. In this case, the DDBs are similarly paired as $\{|\phi_{jk}^{\pm}\rangle,|l_{\mathbb{T}}\rangle:(j,k)\in\mathbb{T}\}$ and $\{|\psi_{jk}^{\pm}\rangle,|l_{\mathbb{T}}\rangle:(j,k)\in\mathbb{T}\}$, where $l_{\mathbb{T}}$ represents the single element in $\{0,1,\dots,d-1\}$ not included in $\mathbb{T}$. For example, with $d=5$ and $\mathbb{T}_{1}=\{(0,1),(2,3)\}$, the value of $l_{\mathbb{T}}$ is 4. A minimum of $d-1$ such partitions is also sufficient.
When $d$ is even, there are $12d-$ 1 DDBs, with the computational basis states $\{|0\rangle,|1\rangle,\cdots,|d-1\rangle\}$ sampled twice. When $d$ is odd, there are $12d-$ 1 DDBs but the computational states $\{|t\rangle:t=0,\cdots,d-1\}$ appear twice. Thus for general dimension $d$, the quantum channel for randomly sampling DDBs is given by:
$2d\times\mathcal{M}(\rho)$ $=2\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})P_{k}+\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\rho P_{jk}^{\pm})\cdot P_{jk}^{\pm}+\text{tr}(\rho Q_{jk}^{\pm})\cdot Q_{jk}^{\pm}\right]$ (19)
$=2\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\sum_{0\leq j<k\leq d-1}[(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)+\rho_{jk}|j\rangle\langle k|+\rho_{kj}|k\rangle\langle j|]$ (20)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\sum_{0\leq j<k\leq d-1}(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)$ (21)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}[\sum_{j=0}^{d-1}\sum_{k=0}^{d-1}(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)-\sum_{k=0}^{d-1}(\rho_{kk}+\rho_{kk})(|k\rangle\langle k|+|k\rangle\langle k|)]$ (22)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}[\sum_{j=0}^{d-1}(d\rho_{jj}|j\rangle\langle j|+\rho_{jj}I+\mathop{\mathrm{tr}}(\rho)|j\rangle\langle j|+\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|)-4\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|]$ (23)
$=\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+\rho+\frac{1}{2}(2d\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|+2\mbox{tr}(\rho)I-4\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|)$ (24)
$=\rho+\mbox{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|.$ (25)
From Eq.(19) to Eq.(20), we use
$\begin{cases}\text{tr}(\rho P_{jk}^{+})=(\rho_{jj}+\rho_{kk}+\rho_{jk}+\rho_{kj})/2,\\ \text{tr}(\rho P_{jk}^{-})=(\rho_{jj}+\rho_{kk}-\rho_{jk}-\rho_{kj})/{2},\\ \text{tr}(\rho Q_{jk}^{+})=(\rho_{jj}+\rho_{kk}+i\rho_{jk}-i\rho_{kj})/{2},\\ \text{tr}(\rho Q_{jk}^{-})=(\rho_{jj}+\rho_{kk}-i\rho_{jk}+i\rho_{kj})/{2},\end{cases}$ (26)
and
$\begin{cases}P_{k}=|k\rangle\langle k|,\\ P_{jk}^{+}=(|j\rangle\langle j|+|k\rangle\langle k|+|j\rangle\langle k|+|k\rangle\langle j|)/2,\\ P_{jk}^{-}=(|j\rangle\langle j|+|k\rangle\langle k|-|j\rangle\langle k|-|k\rangle\langle j|)/2,\\ Q_{jk}^{+}=(|j\rangle\langle j|+|k\rangle\langle k|-i|j\rangle\langle k|+i|k\rangle\langle j|)/2,\\ Q_{jk}^{-}=(|j\rangle\langle j|+|k\rangle\langle k|+i|j\rangle\langle k|-i|k\rangle\langle j|)/2.\end{cases}$ (27)
Denote the following symbols.
$\begin{cases}a_{jk}=\rho_{jj}+\rho_{kk},\\ b_{jk}^{\pm}=\rho_{jk}\pm\rho_{kj},\\ A_{jk}=|j\rangle\langle j|+|k\rangle\langle k|\\ B_{jk}^{\pm}=|j\rangle\langle k|\pm|k\rangle\langle j|.\\ \end{cases}$ (28)
Thus we have
$\begin{cases}\mathop{\mathrm{tr}}(\rho P^{+}_{jk})P^{+}_{jk}=\frac{1}{4}(a_{jk}+b_{jk}^{+})(A_{jk}+B_{jk}^{+})\\ \mathop{\mathrm{tr}}(\rho P^{-}_{jk})P^{-}_{jk}=\frac{1}{4}(a_{jk}-b_{jk}^{+})(A_{jk}-B_{jk}^{+})\\ \mathop{\mathrm{tr}}(\rho Q^{+}_{jk})Q^{+}_{jk}=\frac{1}{4}(a_{jk}+ib_{jk}^{-})(A_{jk}-iB_{jk}^{-})\\ \mathop{\mathrm{tr}}(\rho Q^{-}_{jk})Q^{-}_{jk}=\frac{1}{4}(a_{jk}-ib_{jk}^{-})(A_{jk}+iB_{jk}^{-}).\\ \end{cases}$ (29)
It is easy to verify that the summation above is equal to $a_{jk}A_{jk}+(b_{jk}^{+}B^{+}_{jk}+b_{jk}^{-}B^{-}_{jk})/2=(\rho_{jj}+\rho_{kk})(|j\rangle\langle j|+|k\rangle\langle k|)+\rho_{jk}|j\rangle\langle k|+\rho_{kj}|k\rangle\langle j|$.
Thus the quantum channel is given by
$\mathcal{M}(\rho)=\frac{1}{2d}[\rho+\mbox{tr}(\rho)I+(d-1)\sum_{k=0}^{d-1}\rho_{kk}|k\rangle\langle k|].$ (30)
As a comparison, when we use the uniform sampling of Clifford measurements or MUB measurements, the quantum channel is given by $\mathcal{M}(\rho)=\frac{1}{d+1}(\rho+\mathop{\mathrm{tr}}(\rho)I)$, where $d=2^{n}$.
For the channel corresponding to uniform sampling from Cliffords and MUBs, the mapping of the matrix elements of a density matrix $\rho$ is as follows:
  • The off-diagonal elements $\rho_{jk}$ (where $j\neq k$) are mapped to $\rho_{jk}/(d+1)$.
  • The diagonal elements $\rho_{jj}$ are mapped to $\frac{\rho_{jj}+\mathrm{tr}(\rho)}{d+1}$.
In contrast, for the channel corresponding to uniform sampling from DDBs, the mapping is:
  • The off-diagonal elements $\rho_{jk}$ (where $j\neq k$) are mapped to $\rho_{jk}/(2d)$.
  • The diagonal elements $\rho_{jj}$ are mapped to $\frac{d\times\rho_{jj}+\mathrm{tr}(\rho)}{2d}$.
The inverse reconstruction channel of uniform sampling DDBs is given by
$\mathcal{M}^{-1}(\rho)$ $=2d\Bigg{[}\rho-\frac{d-1}{d}\sum_{k=0}^{d-1}\mbox{tr}(\rho P_{k})P_{k}\Bigg{]}-\frac{\mbox{tr}(\rho)}{d}I.$ (31)
It also includes the linear combination of $\rho$ and $I$, with additional corrections to the diagonal terms.
Appendix B Efficient random projections in DDB-ST versus Clifford-ST
On near-term quantum devices, implementing random Clifford circuits using their standard decomposition—typically involving $O(n^{2})$ elementary gates—remains experimentally challenging for large $n$. A commonly used workaround on optical quantum platforms is to project onto random stabilizer states, which is mathematically equivalent to applying a random Clifford unitary followed by a computational basis measurement.
However, directly sampling $n$-qubit stabilizer states is computationally intensive, as the total number of such states is $O(2^{n^{2}})$. A standard approach is to generate $n$ independent stabilizer generators $\{g_{i}\}_{i=1}^{n}$ and compute the corresponding projector as:
$|\psi\rangle\langle\psi|=\frac{1}{2^{n}}\prod_{i=1}^{n}(I+g_{i}).$
The resulting stabilizer state vector takes the form $|\psi\rangle=\sum_{x\in\{0,1\}^{n}}a_{x}|x\rangle$, which generally contains $2^{n}$ amplitudes. A naive construction based on this expression requires more than $O(2^{3n})$ operations.
To improve efficiency, an alternative method is to directly compute the stabilizer state using its canonical expression:
$|\Psi\rangle=\frac{1}{\sqrt{2^{r}}}\sum_{k\in A}\omega^{q(k)}|k\rangle,$ (32)
where $\omega^{q(k)}$ is a phase factor taking values in $\{1,-1,i,-i\}$, and
$A=\{Ru+t\mid u\in\mathbb{Z}_{2}^{r}\},$
with $R$ being a nonsingular $n\times r$ binary matrix, and $t\in\mathbb{Z}_{2}^{N}$ a fixed vector.
Since the number of amplitudes to be evaluated is $2^{n}$, the overall complexity of constructing such a state remains $O(2^{n}n^{3})$, even under optimized pre-processing techniques struchalin2021experimental .
By contrast, the total number of DDB states in a $d$-dimensional system is $2d^{2}-d$, and sampling from these DDB states is significantly more straightforward. With the sampling stategy, the computational basis state $|j\rangle$ is sampled with probability $2/(2d^{2})$, while the other nontrivial DDB states is sampled with probability $1/d^{2}$.
To avoid implementing the DDB circuits, the equivalent method is to perform projective measurement onto a randomly selected DDB state.
First, we select a computational basis probability $1/d$. Each state $|j\rangle$ (where $j=0,\dots,d-1$) is then sampled with probability $1/d$. Thus $|j\rangle$ is selected with probability $1/d^{2}$.
Next, with probability $1-1/d$, we select a non-computational basis. In this case, two distinct integers $m,n\in\{0,\dots,d-1\}$ are selected such that $10\leq m<n\leq d-$ 1. Then, one of the following four superpositions is chosen randomly with equal probability ($1/4$ each):
$\frac{1}{\sqrt{2}}\left(|m\rangle+|n\rangle\right),\quad\frac{1}{\sqrt{2}}\left(|m\rangle-|n\rangle\right),\quad\frac{1}{\sqrt{2}}\left(|m\rangle+i|n\rangle\right),\quad\frac{1}{\sqrt{2}}\left(|m\rangle-i|n\rangle\right).$
Thus, the probability of selecting any nontrivial DDB state is:
$\frac{1-1/d}{\binom{d}{2}}\times\frac{1}{4}=\frac{1}{2d^{2}}.$
In this process, each DDB state is selected according to the designed probability distribution. The computational resources required are only $O(1)$. In contrast to Clifford-ST, where the projection state is determined through computationally intensive methods, the preprocessing overhead in DDB-ST is exponentially smaller. This is primarily because only two nonzero amplitudes need to be determined, while all other components are zero, eliminating the need for further computation.
Appendix C Proof of proposition 1: Constant-time post-processing in a single measurement for DDB-ST
One of the most significant advantages of DDB-ST is that, for any observable to be estimated, the classical post-processing per single-shot measurement has constant complexity, involving at most four arithmetic operations when $\operatorname{tr}(O)/d$ is deferred to the final step.
Proof. We evaluate the post-processing formula
$\operatorname{tr}\left[\mathcal{M}^{-1}(U_{k}^{\dagger}|j\rangle\langle j|U_{k})\cdot O\right],$
where $\mathcal{M}^{-1}$ is the inverse channel defined by Eq. (31) with $P_{k}=|k\rangle\langle k|$ being projectors onto the computational basis. Let $\rho=U_{k}^{\dagger}|j\rangle\langle j|U_{k}=|\psi\rangle\langle\psi|$, where $|\psi\rangle\in\mathcal{S}_{\text{DDB}}$ is the snapshot vector. We express the observable as $O=\sum_{m,n=0}^{d-1}O_{mn}|m\rangle\langle n|$.
There are three cases:
(1) Computational basis state:
$|\psi\rangle=|t\rangle\quad\Rightarrow\quad\rho=|t\rangle\langle t|.$
Then:
$\operatorname{tr}(\rho P_{k})=\delta_{tk},\quad\operatorname{tr}(\rho)=1,$
so
$\sum_{k=0}^{d-1}\operatorname{tr}(\rho P_{k})P_{k}=P_{t}=|t\rangle\langle t|,$
and thus:
$\mathcal{M}^{-1}(\rho)=2d\left[|t\rangle\langle t|-\frac{d-1}{d}|t\rangle\langle t|\right]-\frac{1}{d}I=2|t\rangle\langle t|-\frac{1}{d}I.$ (33)
Hence:
$\operatorname{tr}[\mathcal{M}^{-1}(\rho)\cdot O]=2O_{tt}-\frac{\operatorname{tr}(O)}{d}.$
(2) Real superposition state:
$|\psi\rangle=\frac{1}{\sqrt{2}}(|m\rangle+|n\rangle),\quad m\neq n.$
Then:
$\rho=|\psi\rangle\langle\psi|=\frac{1}{2}\left(|m\rangle\langle m|+|n\rangle\langle n|+|m\rangle\langle n|+|n\rangle\langle m|\right),$
and
$\operatorname{tr}(\rho P_{k})=\begin{cases}\frac{1}{2},&k=m\text{ or }n,\\ 0,&\text{otherwise}.\end{cases}$
So:
$\sum_{k}\operatorname{tr}(\rho P_{k})P_{k}=\frac{1}{2}(P_{m}+P_{n}),$
and
$\mathcal{M}^{-1}(\rho)=2d\left[\rho-\frac{d-1}{2d}(P_{m}+P_{n})\right]-\frac{1}{d}I=P_{m}+P_{n}+d(|m\rangle\langle m|+|n\rangle\langle n|)-\frac{1}{d}I.$ (34)
With the expression of $O$, we have
$\operatorname{tr}[\mathcal{M}^{-1}(\rho)\cdot O]=O_{mm}+O_{nn}+2d\,\text{Re}(O_{mn})-\frac{\operatorname{tr}(O)}{d}.$
Here we use $O_{mn}=O^{*}_{nm}$, and $O_{mn}=\text{Re}(O_{mn})+i\text{Im}(O_{mn})$.
(3) Imaginary superposition state:
$|\psi\rangle=\frac{1}{\sqrt{2}}(|m\rangle+i|n\rangle).$
Then:
$\rho=\frac{1}{2}(|m\rangle\langle m|+|n\rangle\langle n|-i|m\rangle\langle n|+i|n\rangle\langle m|),$
and again
$\sum_{k}\operatorname{tr}(\rho P_{k})P_{k}=\frac{1}{2}(P_{m}+P_{n}).$
So:
$\mathcal{M}^{-1}(\rho)=2d\left[\rho-\frac{d-1}{2d}(P_{m}+P_{n})\right]-\frac{1}{d}I=P_{m}+P_{n}+d(i|n\rangle\langle m|-i|m\rangle\langle n|)-\frac{1}{d}I.$ (35)
Similarly, we have
$\operatorname{tr}[\mathcal{M}^{-1}(\rho)\cdot O]=O_{mm}+O_{nn}-2d\,\text{Im}(O_{mn})-\frac{\operatorname{tr}(O)}{d}.$
Conclusion: Each of the above formulas involves only a constant number of entries from $O=[O_{mn}]$, and requires a constant number of arithmetic operations. Hence, the classical computational cost of each post-processing step is independent of $d$, i.e., $O(1)$. Specifically, the calculation of $\frac{\operatorname{tr}(O)}{d}$ can be incorporated into the final averaging step, so each post-processing requires at most four arithmetic operations. ∎
Appendix D Proof of theoreom 2: performance guarantee
The predicted observable is $O$. Denote its traceless part as $O_{0}=O-\mbox{tr}(O)I/d$. If we uniformly sample the DDBs for state $\sigma$ as introduced above, the sampling complexity is linearly dependent on the variance
$\mathbb{E}_{U\sim\mathcal{U}}\sum_{b\in\{0,1\}^{n}}\langle b|U\sigma U^{\dagger}|b\rangle\cdot\langle b|U\mathcal{M}^{-1}(O_{0})U^{\dagger}|b\rangle^{2}.$ (36)
We have
$\mathcal{M}^{-1}(O_{0})=2d\left[O_{0}-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(O_{0}P_{k})P_{k}\right].$ (37)
Denote $\mathcal{M}^{-1}(O_{0})=2d\times o$, where $o=O_{0}-\frac{d-1}{d}\sum_{k=0}^{d-1}\text{tr}(O_{0}P_{k})P_{k}$.
By the definition of $o$, we can calculate the relationship between its matrix elements and those of the matrix $O_{0}$:
$&\text{tr}(oP_{k})=\frac{1}{d}\text{tr}(O_{0}P_{k})\\ &\text{tr}(o|j\rangle\langle k|)=\text{tr}(O_{0}|j\rangle\langle k|).\\ \end{aligned}\right.$ (38)
This means that the diagonal elements of the operator $o$ correspond to the reciprocal of the diagonal elements of the operator $O_{0}$ with a denominator of $d$. The operators $o$ and $O_{0}$ share the same non-diagonal elements. The variance for unknown state $\sigma$ is then expressed as follows:
$\lVert O_{0}\rVert_{\sigma}^{2}$ $=\sum_{k=0}^{d-1}\frac{2\text{tr}(\sigma P_{k})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P_{k})+\sum_{0\leq j<k\leq d-1}[\frac{\text{tr}(\sigma P^{+}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P^{+}_{jk})+\frac{\text{tr}(\sigma Q^{+}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})Q^{+}_{jk})]$ (39)
$+\frac{\text{tr}(\sigma P^{-}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})P^{-}_{jk})+\frac{\text{tr}(\sigma Q^{-}_{jk})}{2d}\cdot\text{tr}^{2}(\mathcal{M}^{-1}(O_{0})Q^{-}_{jk})]$
$=4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\sigma P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\sigma Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right].$
D.1 Upper bound for the worst case
Since $\sigma$, $P_{k}$, $P_{jk}^{\pm}$, and $Q_{jk}^{\pm}$ are all quantum states, their inner products satisfy $0\leq\operatorname{tr}(\sigma P_{k}),\operatorname{tr}(\sigma P_{jk}^{\pm}),\operatorname{tr}(\sigma Q_{jk}^{\pm})\leq 1$. By upper bounding each of these terms by 1, we obtain the following bound on the variance:
$\lVert O_{0}\rVert_{\sigma}^{2}\leq 4d\sum_{k=0}^{d-1}\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}^{2}(oQ_{jk}^{\pm})\right]).$ (40)
Denote the matrix elements of $o$ as $o_{jk}$, where $j,k\in\{0,\cdots,d-1\}$. We rewrite Eq. (40) as $\lVert O_{0}\rVert_{\sigma}^{2}\leq 2d\times T$, where
$T\doteq$ $\sum_{k=0}^{d-1}2\text{tr}^{2}(oP_{k})+\sum_{0\leq j<k\leq d-1}\left[\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}^{2}(oQ_{jk}^{\pm})\right]$
$=$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\frac{1}{4}\sum_{0\leq j<k\leq d-1}\big{[}(o_{jj}+o_{kk}+o_{jk}+o_{kj})^{2}+(o_{jj}+o_{kk}-o_{jk}-o_{kj})^{2}$
$+(o_{jj}+o_{kk}+io_{jk}-io_{kj})^{2}+(o_{jj}+o_{kk}-io_{jk}+io_{kj})^{2}\big{]}$
$=$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}\big{[}(o_{jj}+o_{kk})^{2}+2o_{jk}o_{kj}\big{]}$
$\leq$ $\sum_{k=0}^{d-1}2o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}\big{[}2(o_{jj}^{2}+o_{kk}^{2})+2o_{jk}o_{kj}\big{]}$
$=$ $2d\sum_{k=0}^{d-1}o_{kk}^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}$
$=$ $\frac{2d}{d^{2}}\sum_{k=0}^{d-1}[d\times o_{kk}]^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}.$
When $d\geq 2$, we have $\frac{2d}{d^{2}}\leq 1$. So, we can deduce the upper bound of $T$.
$T\leq\sum_{k=0}^{d-1}[d\times o_{kk}]^{2}+\sum_{0\leq j<k\leq d-1}2o_{jk}o_{kj}=\text{tr}(O_{0}^{2}).$
Here we use the relation in Eq. (38).
Thus the upper bound of $\lVert O_{0}\rVert_{\sigma}^{2}$ can be deduced. For each unknown state $\sigma$, the expectation values of $\text{tr}(\sigma P_{k})$, $\text{tr}(\sigma P_{jk}^{\pm})$, and $\text{tr}(\sigma Q_{jk}^{\pm})$ are no bigger than 1. Then we have
$\lVert O_{0}\rVert_{\text{ shadow}}^{2}=\max_{\sigma:\text{ state }}\lVert O_{0}\rVert_{\sigma}^{2}\leq 2dT\leq 2d\times\text{tr}(O_{0}^{2}).$ (41)
Consider the variance in Eq. (39), the first part $4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})$ tends to zero as $d$ increases. As we have $\text{tr}^{2}(oP_{k})=\frac{\mathop{\mathrm{tr}}^{2}(O_{0}|k\rangle\langle k|)}{d^{2}}\leq\mathop{\mathrm{tr}}(O^{2}_{0})/d^{2}$ for all $k$. Thus
$4d\sum_{k=0}^{d-1}\text{tr}(\sigma P_{k})\cdot\text{tr}^{2}(oP_{k})\leq\frac{4\mathop{\mathrm{tr}}(O^{2}_{0})}{d}.$ (42)
Then for BN observable, the efficiency of sampling complexity is just related to the following part:
$V_{\mbox{diag}}=2d\times\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\sigma P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\sigma Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right].$ (43)
One worst case could happen when the unknown state and the observable are the same as one of the nontrivial DDB states. For example, $\sigma=P_{01}^{+}$ and $O=P_{01}^{+}$. Then $O_{0}=P_{01}^{+}-I/d$, $o=\frac{d-2}{2d^{2}}(|0\rangle\langle 0|+|1\rangle\langle 1|)+\frac{|0\rangle\langle 1|+|1\rangle\langle 0|}{2}$. Thus $\mathop{\mathrm{tr}}(oP_{jk})>1/2$ and $V_{\mbox{diag}}>d$. Then in this case, the sampling complexity is linear dependent with $d$.
D.2 Average performance analysis
If we sample $|\phi\rangle=U|0\rangle$ with Haar measure, the average state will be
$\int_{U(2^{n})}U|0\rangle\langle 0|U^{{\dagger}}\mathrm{d}\mu(U)=I/2^{n}\;.$ (44)
Thus the variance of state $\sigma=I/d$ exhibits the average performance when the output $\rho$ is randomly and uniformly generated.
When the unknown state is $\sigma=I/d$, we have
$\mathop{\mathrm{tr}}(\sigma P_{k})=\mathop{\mathrm{tr}}(\sigma P^{\pm}_{jk})=\mathop{\mathrm{tr}}(\sigma Q^{\pm}_{jk})=1/d.$ (45)
Take Eq. (45) into Eq. (39), we have
$\lVert O_{0}\rVert_{I/d}^{2}\leq\frac{1}{d}2dT\leq 2\text{tr}(O_{0}^{2}).$ (46)
Thus, the average performance is efficient for bounded-norm observables. This value is approximately twice that of the average performance obtained by uniformly sampling from a complete set of mutually unbiased bases (MUBs), which yields $(1+1/2^{n})\operatorname{tr}(O_{0}^{2})$ for $d=2^{n}$. However, a complete set of $d+1$ MUBs is known to exist only in prime power dimensions.
Appendix E Proof of lemma 1: approximate average state case
If the state $\rho$ is approximately DDB-average, then its deviation from the completely mixed state $I/d$ is small. Specifically, we have
$\left|\text{tr}(\rho|\phi\rangle\langle\phi|)-\frac{1}{d}\right|\leq\frac{O(\text{poly}(\log d))}{d}$ (47)
for all snapshots $|\phi\rangle\in\mathcal{S}_{\text{DDB}}$. This implies that $\text{tr}(\rho|\phi\rangle\langle\phi|)\leq\frac{O(\text{poly}(\log d))+1}{d}$. Substituting this into Eq. (39), we can deduce
$\lVert O_{0}\rVert_{\rho}^{2}$ $=4d\sum_{k=0}^{d-1}\text{tr}(\rho P_{k})\cdot\text{tr}^{2}(oP_{k})+2d\sum_{0\leq j<k\leq d-1}\left[\text{tr}(\rho P_{jk}^{\pm})\cdot\text{tr}^{2}(oP_{jk}^{\pm})+\text{tr}(\rho Q_{jk}^{\pm})\cdot\text{tr}^{2}(oQ_{jk}^{\pm})\right]$ (48)
$\leq O(\text{poly}(\log d)+1)\times 2T$
$\leq O(\text{poly}(\log d))\times\text{tr}(O^{2}_{0}).$
Denote $\epsilon_{1}=\frac{O(\text{poly}(\log d))}{d}$. It is an interesting question to characterize the proportion of randomly chosen states for various levels of deviation $\epsilon_{1}$, such as $\frac{\log d}{d}$, $\frac{\log^{2}d}{d}$, $\frac{\log^{3}d}{d}$, and so forth.
E.1 More exact variance formula
We now present a more concise expression for $V_{\text{diag}}$ in Eq. (43) by substituting the form given in Eq. (26). The variance for $\rho$ and a BN-observable $O$, given by $\lVert O_{0}\rVert_{\sigma}^{2}$, can be expressed as $V_{\text{diag}}$ plus a term that vanishes as $d\to\infty$. Hence, the growth of $\lVert O_{0}\rVert_{\sigma}^{2}$ is determined by the behavior of $V_{\text{diag}}$. If $V_{\text{diag}}$ scales polynomially, then $\lVert O_{0}\rVert_{\sigma}^{2}$ also exhibits polynomial scaling; conversely, if $V_{\text{diag}}$ scales exponentially, $\lVert O_{0}\rVert_{\sigma}^{2}$ will follow an exponential growth pattern as well.
$V_{\text{diag}}$ $=2d\sum_{0\leq j<k\leq d-1}\left[\operatorname{tr}(\sigma P_{jk}^{\pm})\cdot\operatorname{tr}^{2}(oP_{jk}^{\pm})+\operatorname{tr}(\sigma Q_{jk}^{\pm})\cdot\operatorname{tr}^{2}(oQ_{jk}^{\pm})\right]$ (49)
$=\frac{d}{4}\sum_{0\leq j<k\leq d-1}\Big{[}(\sigma_{jj}+\sigma_{kk}+\sigma_{jk}+\sigma_{kj})(o_{jj}+o_{kk}+o_{jk}+o_{kj})^{2}$
$\quad+(\sigma_{jj}+\sigma_{kk}-\sigma_{jk}-\sigma_{kj})(o_{jj}+o_{kk}-o_{jk}-o_{kj})^{2}$
$\quad+(\sigma_{jj}+\sigma_{kk}+i\sigma_{jk}-i\sigma_{kj})(o_{jj}+o_{kk}+io_{jk}-io_{kj})^{2}$
$\quad+(\sigma_{jj}+\sigma_{kk}-i\sigma_{jk}+i\sigma_{kj})(o_{jj}+o_{kk}-io_{jk}+io_{kj})^{2}\Big{]}$
$=\frac{d}{4}\sum_{0\leq j<k\leq d-1}\Big{[}4(\sigma_{jj}+\sigma_{kk})(o_{jj}+o_{kk})^{2}$
$\quad+4(\sigma_{jk}+\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}+o_{kj})$
$\quad-4(\sigma_{jk}-\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}-o_{kj})$
$\quad+8(\sigma_{jj}+\sigma_{kk})o_{jk}o_{kj}\Big{]}$
$=d\sum_{0\leq j<k\leq d-1}[(\sigma_{jj}+\sigma_{kk})(o_{jj}+o_{kk})^{2}+(\sigma_{jk}+\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}+o_{kj})-(\sigma_{jk}-\sigma_{kj})(o_{jj}+o_{kk})(o_{jk}-o_{kj})$
$\quad+2(\sigma_{jj}+\sigma_{kk})o_{jk}o_{kj}].$
Now, consider the case where $O$ is an off-diagonal observable, meaning $O_{jj}=o_{jj}=0$ for all $j=0,\ldots,d-1$. Under this condition, the expression simplifies to:
$V_{\text{diag}}=2d\sum_{0\leq j<k\leq d-1}(\sigma_{jj}+\sigma_{kk})o_{jk}o_{kj}.$
Appendix F Efficient estimation of stabilizer states with DDB-ST
Property 2 in the main text is easily verified through direct calculation.
Lemma 2.
Given arbitrary $n$-qubit stabilizer state $|\Psi\rangle$ in Eq. (32), we can efficiently construct a Clifford circuit $T$, composed of elementary gates from the set $\{S,\text{CZ},\text{CX}\}$, such that $|\Phi\rangle=T|\Psi\rangle=|0\rangle^{\otimes(n-r)}\otimes|\Phi_{r}\rangle$. Here, CZ and CX denote the controlled-Z and controlled-X gates, respectively, and $|\Phi_{r}\rangle$ is an $r$-qubit stabilizer state with $2^{r}$ nonzero amplitudes.
Proof: By Eq. (42) of bravyi2019simulation , any stabilizer state can be expressed as
$|\Psi\rangle=\omega U_{C}U_{H}|s\rangle,$
where $U_{C}$ and $U_{H}$ are $C$-type and $H$-type Clifford operators, $s\in\{0,1\}^{n}$ is a basis vector, and $\omega$ is a complex phase factor. The $C$-type Clifford operators are those that can be decomposed into gates from the set $\{S,\text{CZ},\text{CX}\}$, while the $H$-type Clifford operators only consist of Hadamard gates. Furthermore, a polynomial-time algorithm is provided to efficiently obtain this CH-form of the stabilizer state using the stabilizer tableaux representation.
Consequently, the target Clifford circuit $T$ can be decomposed into $U_{C}^{\dagger}$, followed by a sequence of SWAP operations (each equivalent to three CX gates). ∎
F.1 Proof of Theorem 3
Now we give the proof of Theorem 3, as stated in the main text.
Proof.
We categorize the cases based on the number of non-zero coefficients present in $|\Psi\rangle\langle\Psi|$ by Eq. (32).
We categorize the cases based on the number of non-zero coefficients present in $|\Psi\rangle\langle\Psi|$ by Eq. (32).
Case 1: $n-\log(\operatorname{poly}(n))\leq r\leq n$. The stabilizer state $|\Psi\rangle$ is approximately DDB-average. Thus, $O(\operatorname{poly}(n))$ samples and computational resources are sufficient for efficient estimation using $n$-qubit DDB-ST (Lemma 1).
Case 2: $0\leq r\leq\log(\operatorname{poly}(n))$. The matrix $|\Psi\rangle\langle\Psi|$ is sparse, containing at most $4^{r}$ non-zero coefficients, making direct computation of $\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ efficient.
Now we consider the general $r$.
State transformation. For a general $r$ and a given stabilizer state $|\Psi\rangle$, we can efficiently construct a Clifford circuit $U$ such that $|\Psi\rangle=U|j\rangle$, where $|j\rangle$ is a computational basis state and $U$ is in CH-form bravyi2019simulation .
We first apply the Gottesman-Knill theorem to decompose $|\Psi\rangle=V|j\rangle$ using standard Clifford generators (Hadamard, Phase, and CNOT gates). This decomposition can be computed in $O(n^{3})$ time with at most $O(n^{2})$ gates. Following the CH-form transformation technique introduced in bravyi2019simulation , each Clifford gate is then translated into CH-form with runtime $O(n)$ per $S$, CZ, and CNOT gate, and $O(n^{2})$ per Hadamard gate. In total, the construction of the CH-form representation $|\Psi\rangle=U|j\rangle$ requires at most $O(n^{5})$ time.
By Lemma S1, we can further identify a unitary operator $T$, composed entirely of SWAP gates, such that
$|\Phi\rangle=T|\Psi\rangle=TU|j\rangle,$
where
$|\Phi\rangle$ $=|0\rangle^{\otimes(n-r)}\otimes\frac{1}{\sqrt{2^{r}}}\sum_{k\in\{0,1\}^{r}}\omega^{q^{\prime}(k)}|k\rangle$ (50)
$=|0\rangle^{\otimes(n-r)}\otimes|\Phi_{r}\rangle.$
The role of $T$ is to permute qubits such that all Hadamard gates in the CH-form circuit are relocated to act on the last $r$ qubits. This transformation can be implemented using at most $O(n)$ SWAP operations and can be found in $O(n)$ time. In other words, we obtain a new stabilizer state $|\Phi\rangle$ where only the first $2^{r}$ components are non-zero.
Reduction to $r$-qubit DDB-ST. The target expectation value can be rewritten as
$\mathrm{tr}(|\Psi\rangle\langle\Psi|O)$ $=\mathrm{tr}(T|\Psi\rangle\langle\Psi|T^{\dagger}\cdot TOT^{\dagger})$ (51)
$=\mathrm{tr}(|\Phi\rangle\langle\Phi|\cdot TOT^{\dagger}),$
where $|\Phi\rangle=|0\rangle^{\otimes(n-r)}\otimes|\Phi_{r}\rangle$, and $|\Phi_{r}\rangle$ is an $r$-qubit DDB-average stabilizer state (by Property 2).
Since the matrix form of $|\Phi\rangle\langle\Phi|$ is block-diagonal with a single nonzero block in the upper-left corner, we have:
$|\Phi\rangle\langle\Phi|=\begin{bmatrix}|\Phi_{r}\rangle\langle\Phi_{r}|&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix},$
and hence only the leading $2^{r}\times 2^{r}$ submatrix of $TOT^{\dagger}$, denoted by $[TOT^{\dagger}]_{2^{r}\times 2^{r}}$, contributes to the trace. Therefore,
$\mathrm{tr}(|\Psi\rangle\langle\Psi|O)=\mathrm{tr}(|\Phi_{r}\rangle\langle\Phi_{r}|\cdot[TOT^{\dagger}]_{2^{r}\times 2^{r}}).$
Since unitary conjugation preserves the Hilbert-Schmidt norm, we have $\operatorname{tr}[(TOT^{\dagger})^{2}]=\operatorname{tr}(O^{2})$, which implies
$\operatorname{tr}([TOT^{\dagger}]_{2^{r}\times 2^{r}}^{2})\leq\operatorname{tr}(O^{2}).$
We can apply $r$-qubit DDB-ST to estimate this quantity efficiently with sample complexity bounded by $O(1)\cdot\operatorname{tr}(O^{2})$. The sample complexity is upper bounded by $O(1)\cdot\mathrm{tr}(O^{2})$, where the constant factor arises from the reduction of the numerator in Eq. (47) from $O(\mathrm{poly}(\log d))$ to 1.
Post-processing in $r$-qubit DDB-ST. In the post-processing step, we evaluate $O(1)\cdot\mathrm{tr}(O^{2})$ expectations of the form
$\mathrm{tr}(\mathcal{M}^{-1}(|\phi_{r}\rangle\langle\phi_{r}|)\cdot[TOT^{\dagger}]_{2^{r}\times 2^{r}}),$
where each $|\phi_{r}\rangle$ is an $r$-qubit DDB state, and $\mathcal{M}^{-1}$ is the inverse channel defined in Eq. (31), with $d=2^{r}$. Let $\tau_{r}:=\mathcal{M}^{-1}(|\phi_{r}\rangle\langle\phi_{r}|)$. By Eqs. (33), (34), and (35), we have
$\tau_{r}=\tau_{r}^{\prime}-\frac{I_{r}}{2^{r}},$ (52)
where $\tau_{r}^{\prime}$ is a $2^{r}$-dimensional operator with at most four nonzero components, and $I_{r}$ is the identity on $\mathbb{C}^{2^{r}}$.
Although $[TOT^{\dagger}]_{2^{r}\times 2^{r}}$ is a submatrix of $TOT^{\dagger}$, its explicit form is not required. Define
$\tau:=(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r},$
then we obtain
$\mathrm{tr}(\tau_{r}\cdot[TOT^{\dagger}]_{2^{r}\times 2^{r}})$ $=\mathrm{tr}(\tau\cdot TOT^{\dagger})=\mathrm{tr}(T^{\dagger}\tau T\cdot O)$ (53)
$=\mathrm{tr}\big{(}T^{\dagger}\big{[}(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}\big{]}T\cdot O\big{)}$
$=\mathrm{tr}\big{(}T^{\dagger}\big{[}(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}^{\prime}\big{]}T\cdot O\big{)}-\frac{1}{2^{r}}\cdot\mathrm{tr}\big{(}T^{\dagger}\big{[}(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}\big{]}T\cdot O\big{)}.$
The first term in the subtraction of Eq. (53) is defined as
$L_{1}:=\mathrm{tr}\big{(}T^{\dagger}\big{[}(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}^{\prime}\big{]}T\cdot O\big{)}.$ (54)
The right-hand side of the subtraction in Eq. (53) is defined as
$L_{2}:=\frac{1}{2^{r}}\,\mathrm{tr}\left(T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}\right]T\cdot O\right).$ (55)
In post-processing, we only calculate the value of $L_{1}$ in Eq. (54) and neglect the value of $L_{2}$ in Eq. (55).
Time complexity to calculate $L_{1}$ in Eq. (54). Since $\tau_{r}^{\prime}$ contains at most four nonzero components, the tensor product $(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}^{\prime}$ also has at most four nonzero entries. Recall that $T$ is composed of at most $O(n)$ CNOT gates and hence acts as a computational basis permutation. Consequently, the operator
$T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}^{\prime}\right]T$
can be computed in $O(n)$ time and remains a sparse matrix with at most four nonzero components.
Therefore, once $T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}^{\prime}\right]T$ is obtained, the final trace
$\mathrm{tr}\left(T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes\tau_{r}^{\prime}\right]T\cdot O\right)$
can be evaluated in constant time, due to the sparsity of the operator involved.
Thus, the total calculation of $L_{1}$ in Eq. (54) can be evaluated with classical time complexity
$O(n^{5})\cdot\left(O(1)\cdot\mathrm{tr}(O^{2})\cdot\frac{\log(1/\sigma)}{\epsilon^{2}}\right)\cdot O(n)\cdot\text{constant}.$ (56)
Here, the term $O(n^{5})$ accounts for the preparation of the state $|\Phi_{r}\rangle$; the term $O(1)\cdot\mathrm{tr}(O^{2})\cdot\frac{\log(1/\sigma)}{\epsilon^{2}}$ corresponds to the sample complexity of $|\Phi_{r}\rangle$ in the $r$-qubit DDB-ST protocol; and $O(n)\cdot\text{constant}$ represents the classical cost of each post-processing step due to the sparsity of the involved operators. The $\epsilon$ is the error for accuracy and $1-\sigma$ is the success probability for confidence level.
Error by neglecting $L_{2}$ in Eq. (55). The value of $L_{2}$ in Eq. (55) can be simplified to $\frac{1}{2^{r}}\sum_{i\in A}O_{ii}$, where $A\subset\{0,1,\dots,2^{n}-1\}$ is an index set of size $|A|=2^{r}$ corresponding to the diagonal positions labeled by the support of
$T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}\right]T.$
Note that $(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}$ is a diagonal projector with exactly $2^{r}$ ones on the diagonal and zeros elsewhere. Since $T$ is a qubit permutation composed of $O(n)$ SWAP or CNOT gates, and the conjugated operator $T^{{\dagger}}$ is equal to $T$. $T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}\right]T.$ is then also a diagonal operation with $2^{r}$ ones determined by $T$ adn zeros at other places.
Thus $L_{2}$ in Eq. (55) is equal to $\frac{1}{2^{r}}$ times the sum of $2^{r}$ diagonal elements of the $2^{n}\times 2^{n}$ matrix $O$, where the selected positions are determined by the support of the permuted projector $T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}\right]T$.
In the general case, the second term can be bounded by
$L_{2}\leq\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}.$ (57)
This bound can be shown as follows. Without loss of generality, assume that the nonzero positions of the diagonal projector $T^{\dagger}\left[(|0\rangle\langle 0|)^{\otimes(n-r)}\otimes I_{r}\right]T$ correspond to the first $2^{r}$ diagonal entries of $O$. Then we can write
$L_{2}=\frac{1}{2^{r}}\sum_{i=1}^{2^{r}}O_{ii}.$
By the Cauchy–Schwarz inequality, we have
$\sum_{i=1}^{2^{r}}O_{ii}\leq\sqrt{2^{r}}\cdot\left(\sum_{i=1}^{2^{r}}O_{ii}^{2}\right)^{1/2}\leq\sqrt{2^{r}}\cdot\left(\sum_{i=1}^{2^{n}}O_{ii}^{2}\right)^{1/2}\leq\sqrt{2^{r}}\cdot\sqrt{\operatorname{tr}(O^{2})}.$
Dividing both sides by $2^{r}$, we obtain the relation in Eq. (57).
In the general case with arbitrary $r$, the estimation error is bounded by
$\epsilon+\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}.$
with the time complexity in Eq. (56).
The term $\epsilon$ accounts for the estimation error introduced by the DDB-ST procedure. The term $\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}$ arises from neglecting the contribution of $L_{2}$ in each estimation step.
Impact of the error term $\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}$. Consider the special case where $O$ is traceless. In this case, $L_{2}$ in Eq. (55) is equal to zero because all diagonal entries $O_{ii}$ are zero. Consequently, the total estimation error reduces to $\epsilon$, and is independent of the term $\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}$.
For the general case, we neglect $L_{2}$ in Eq. (55). While for the standard DDB-ST applied to the full $n$-qubit system, the corresponding correction term simplifies to $\operatorname{tr}(O)/2^{n}$, which can be computed directly since $\operatorname{tr}(O)$ is assumed to be known.
When $r$ is small—for example, $0\leq r\leq\log(\mathrm{poly}(n))$—the term $\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}$ can have a significant impact on the overall estimation error. However, in this regime, the state $|\Psi\rangle\langle\Psi|$ is sparse, and the exact computation of $\operatorname{tr}(|\Psi\rangle\langle\Psi|O)$ remains tractable through direct calculation.
In contrast, when $r$ becomes large—for instance, $r=O(n)$—the term $\frac{\operatorname{tr}(O^{2})}{\sqrt{2^{r}}}$ becomes negligible as we have $\operatorname{tr}(O^{2})\leq\operatorname{poly}(n)$ for BN observable $O$. Although exact computation of $\operatorname{tr}(|\Psi\rangle\langle\Psi|O)$ is no longer feasible in this case, the error introduced by neglecting the $L_{2}$ term remains well-controlled due to the exponential suppression from the denominator $2^{r}$. And it is efficient to perform the $r$-qubit DDB-ST on the transformed state.
Finally, we would like to point out that it is not feasible to combine the techniques of Clifford-ST and DDB-ST to efficiently estimate $\operatorname{tr}(\rho O)$ for arbitrary $n$-qubit quantum state $\rho$ and any BN observable $O$ observable with known trace.
The main bottleneck of Clifford-ST lies in its post-processing cost, namely, it does not guarantee that $\operatorname{tr}(U_{k}^{\dagger}|j\rangle\langle j|U_{k}O)$ can always be computed in polynomial time for each single-shot measurement. While $U_{k}^{\dagger}|j\rangle\langle j|U_{k}$ is a stabilizer state and $O$ is a BN observable, which seemingly allows for efficient estimation of $\operatorname{tr}(U_{k}^{\dagger}|j\rangle\langle j|U_{k}O)$ by DDB-ST as a subroutine (Theorem 3), the issue lies in the accumulated error.
Specifically, in the post-processing formula of Clifford-ST (Eq. (7) in the main text), each term $\operatorname{tr}(U_{k}^{\dagger}|j\rangle\langle j|U_{k}O)$ is multiplied by a coefficient of $12^{n}+$ 1. As a result, the DDB-ST subroutine used to estimate $\operatorname{tr}(U_{k}^{\dagger}|j\rangle\langle j|U_{k}O)$ must achieve exponentially small error in order to ensure the overall accuracy of the weighted term $(2^{n}+1)\operatorname{tr}(U_{k}^{\dagger}|j\rangle\langle j|U_{k}O)$. This, in turn, requires an exponential number of samples in DDB-ST, which severely limits the efficiency of the estimation and ultimately negates the potential advantages of such a hybrid approach.